### v2

# Finite Groups of Symplectic Automorphisms of Supersingular K3 surfaces in Odd Characteristics

Abstract.
In 2009, Dolgachev–Keum show that finite groups of tame symplectic automorphisms of K3 surfaces in positive characteristics are subgroups of the Mathieu group of degree $23$ under mild conditions. In this paper, we utilize lattice-theoretic methods to investigate symplectic actions of finite groups $G$ on K3 surfaces in odd characteristics. In the tame cases (i.e., the order of $G$ is coprime with $p$) and all the superspecial cases, one can associate with the action a Leech pair which can be detected via Höhn–Mason’s list. Notice that the superspecial cases have been resolved by Ohashi–Schütt. If the K3 surface is supersingular with Artin invariant at least two, we develop a new machinery called $p$-root pairs to detect possible symplectic finite group actions (without the assumption of tameness). The concept of $p$-root pair is closely related to root systems and Weyl groups. In particular, we recover many results by Dolgachev–Keum with a different method and give an upper bound for the exponent of $p$ in $|G|$.
1. Introduction
K3 surfaces play an important role in many areas of mathematics and the study of their automorphism groups has a long and rich history. Nikulin [Nik79b] first systematically studied automorphisms of K3 surfaces via lattice theory. After that, in his pioneering work [Muk88], Mukai proved (using the so-called Mathieu representations) that a finite group acts faithfully and symplectically on a complex K3 surface if and only if it is a subgroup of the Mathieu group $M_{23}$ with the induced action on the set of $24$ points having at least $5$ orbits. The problem of classifying (symplectic) automorphism groups of K3 surfaces over fields of positive characteristics turns out to be more complicated. In [DK09], Dolgachev and Keum extended Mukai’s methods to classify finite groups $G$ that act symplectically and faithfully on K3 surfaces over an algebraically closed field of characteristic $p>0$ under the tameness condition that $\gcd(|G|,p)=1$. They also showed that if $p\geq 13$, then such $G$ has to be tame. Ohashi–Schütt [OS24] recently completely classified maximal finite groups of symplectic automorphisms of superspecial K3 surfaces in any positive characteristic.
In this paper, we investigate finite groups acting symplectically on K3 surfaces in positive characteristic $p\geq 3$ via the induced action of the group on the Néron–Severi lattice. Our method can be applied to group actions without the tameness assumption. Before stating our main results, we fix some terminologies.
Throughout this paper, let $p$ be an odd prime, and $k$ be any algebraically closed field with $\operatorname{char}(k)=p$. Let $X$ be a K3 surface in characteristic $p$ (which means that it is a K3 surface over $k$ in this paper). We denote the Néron–Severi lattice of $X$ by $\operatorname{NS}(X)$ with rank $\rho(X)$ (which is the Picard number of $X$). It is known that $\operatorname{NS}(X)$ is an even lattice with signature $(1,\rho(X)-1)$. The discriminant group of the lattice $\operatorname{NS}(X)$ is denoted by $A_{\operatorname{NS}(X)}$. We denote the automorphism group of $X$ by $\operatorname{Aut}(X)$. An automorphism $g$ of $X$ is symplectic if and only if $g$ acts identically on $H^{0}(X,\omega_{X})$. Here $\omega_{X}$ is the canonical line bundle on $X$. We denote by $\operatorname{Aut}^{s}(X)$ the group of symplectic automorphisms of $X$. Given a finite subgroup $G\leq\operatorname{Aut}^{s}(X)$, there is an induced action of $G$ on the Néron–Severi lattice $\operatorname{NS}(X)$.
In general, if we have an action of a finite group $G$ on a lattice $L$, we denote by $L^{G}$ the invariant sublattice, and by $S_{G}(L)$ the orthogonal complement of $L^{G}$ in $L$. We call the lattice $S_{G}(L)$ the covariant lattice associated with the group action. We put $S_{G}(X)=S_{G}(\operatorname{NS}(X))$ for short. This lattice is always rootless because any finite group action on $X$ fixes certain ample class (see discussion in §3). The pair $(G,S_{G}(X))$ is a so-called Leech pair, see Definition 2.8. We denote by $\mathbb{L}$ the Leech lattice, and by $\operatorname{Co}_{0}$ the Conway group (see §2.4 for more details.). One key point of our approach is that a Leech pair $(G,S)$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$ if $\operatorname{rank}(S)+\ell(A_{S})\leq 24$, see Proposition 2.9. Here $A_{S}$ represents the discriminant group of $S$, and for a finitely generated abelian group $A$, $\ell(A)$ is the minimal number of generators. Höhn–Mason [HM16] classified all possible sublattices of $\mathbb{L}$ which are covariant lattices for certain finite group actions on $\mathbb{L}$ (there are $290$ cases). This is the main lattice-theoretic input for our approach.
Suppose a group $G$ acts faithfully on a lattice $L$ such that $L^{G}=\{0\}$ and $G$ acts trivially on $A_{L}$. Then we denote by $\widetilde{G}$ the group of all automorphisms of $L$ which act trivially on $A_{L}$. We call $\widetilde{G}$ and $(\widetilde{G},L)$ the enlargement of $G$ and $(G,L)$ respectively.
The K3 surface $X/k$ is said to be (Shioda-)supersingular if $\rho(X)$ reaches the maximal possibility $22$. When $X$ is supersingular,
$A_{\operatorname{NS}(X)}\cong(\mathbb{Z}/p\mathbb{Z})^{2\sigma},1\leq\sigma\leq 10.$
The number $\sigma$ is called the Artin invariant of $X$. For each $1\leq\sigma\leq 10$, the Néron–Severi lattices of all supersingular K3 surfaces with Artin invariant $\sigma$ are isomorphic, and we denote it by $N_{p,\sigma}$, see Proposition-Definition 2.13. A supersingular K3 surface is called superspecial if its Artin invariant is 1. It is known that there exists a unique superspecial K3 surface over $k$ up to isomorphism. If a K3 surface in characteristic $p$ is not supersingular, then it is of finite height with height $1\leq h\leq 10$, see discussions in §4.
Employing the lattice-theoretic method, we will prove the following theorem.
Theorem  1.1.
Suppose $G$ is a finite group acting symplectically and faithfully on a K3 surface $X$ in characteristic $p\geq 3$. Suppose moreover $\gcd(|G|,p)=1$. Then
  1. (1)
    If $X$ is supersingular with Artin invariant $\sigma$, then $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$. Moreover, the enlarged pair $(\widetilde{G},S_{G}(X))$ is isomorphic to a pair in Table LABEL:table:_sublist_of_HM.
  2. (2)
    If $X$ is of finite height $h$, then $\operatorname{rank}(S_{G}(X))\leq 21-2h$, and there exists a primitive embedding of $S_{G}(X)$ into the K3 lattice.
In any case, $(G,S_{G}(X))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$ and $\operatorname{rank}(S_{G}(X))+\ell(A_{S_{G}(X)})\leq 22$.
By Höhn–Mason’s table in [HM16], a Leech pair $(G,S)$ such that $\operatorname{rank}(S)+\ell(A_{S})\leq 22$ implies that $G$ is a subgroup of $M_{23}\leq\operatorname{Co}_{0}$. Therefore, from Theorem 1.1, any finite group of tame symplectic automorphisms of a K3 surface in odd characteristic is a subgroup of $M_{23}$. This is highly compatible with the classification by Dolgachev–Keum [DK09, Theorem 4.7] and the recent result by Ohashi–Schütt [OS24, Proposition 8.2, Theorem 8.4].
Let $X$ be supersingular with Artin invariant $\sigma$. We can extend the action of $G$ on $N_{p,\sigma}$ to $N_{p,1}$. The lattice $S_{G}(N_{p,\sigma})$ is rootless, while $S_{G}(N_{p,1})$ might contain roots. We denote by $R$ the root lattice of $S_{G}(N_{p,1})$. Apparently, $G$ acts on $R$. Let $G_{R}$ be the image of $G$ in $\operatorname{Aut}(R)$, and $G_{L}=\mathrm{Ker}(G\to\operatorname{Aut}(R))$. Then we have exact sequence:
$1\rightarrow G_{L}\rightarrow G\rightarrow G_{R}\rightarrow 1$
Further exploring the action of $G_{R}$ on $R$, we introduce the concept of (pseudo) p-root pairs (see §5). We summarized most of our results on supersingular K3 surfaces proved via $p$-root pairs in §5 and §6 as follows:
Theorem  1.2.
Let $X$ be a supersingular K3 surface over $k$, with Artin invariant $\sigma$. Let $G$ be a finite group acting faithfully and symplectically on $X$. Denote by $W(R)$ the Weyl group of $R$.
  1. (1)
    If $G$ is tame then $G=G_{L}$.
  2. (2)
    $G_{R}\cap W(R)$ is always a $p$-group.
  3. (3)
    If $p\geq 13$, $G$ is always tame.
  4. (4)
    If $p=11$, then the exponent of $p$ in the order of $G$ is at most $1$.
Combining both analysis on $G_{L}$ and $G_{R}$, we also obtain an upper bound (Proposition 6.8) on the exponent of $p$ in $|G|$ for a finite group $G$ of symplectic automorphisms of a K3 surface in characteristic $p=3,5,7$.
Now let us recall how our approach is related with earlier works. Kondō [Kon98] gave another proof of Mukai’s results which is lattice-theoretic and motivates our proof here. Ogus [Ogu79] (respectively Bragg–Lieblich [BL18]) proved the crystalline Torelli theorem for supersingular K3 surfaces in char $p\geq 5$ (respectively $p\geq 3$), which makes the lattice theory a powerful tool in the study of $\operatorname{Aut}^{s}(X)$. Jang [Jan16] studied the action of a symplectic automorphism on the discriminant group of the Néron–Severi group in odd characteristics which enables us to obtain Leech pairs. The key input on lattice-theoretic side is from Höhn–Mason [HM16], Jang [Jan16] and also inspired by methods developed in Laza–Zheng [LZ22].
Structure of the paper: In §2, we collect lattice-theoretic results that will be used in our paper. A key result is Proposition 2.9. In §3 we introduce supersingular K3 surfaces and recall the crystalline Torelli theorem by Ogus, Bragg and Lieblich. In particular, we give the motivation for the concept of $p$-root pairs. In §4, we prove Theorem 1.1 and discuss its relation with previous works. In §5, we developed the new machinery of (pseudo) $p$-root pairs, and in §6 we apply the new machinery to reprove an important result by Dolgachev–Keum (Theorem 1.2, (3)) in supersingular cases and give upper bounds on the wild degree of finite groups of symplectic automophisms of K3 surfaces in characteristics $p=3,5,7,11$. In Appendix A, we include the full list of Höhn–Mason pairs admitting primitive embeddings in $N_{p,1}$.
2. Lattice-Theoretical Results
In this section, we introduce some terminologies and results from lattice theory that will be used. We refer to [Nik79a], [CS99], [LZ22, Appendix A] for more details.
2.1. Integral Lattices and $p$-Adic Lattices
Let $R$ be a PID (principal integral domain) with fractional field $K$. An $R$-lattice is a free $R$-module $L$ of finite rank endowed with a non-degenerate symmetric bilinear form $\phi:L\times L\rightarrow R$, which can be extended to a non-degenerate symmetric $K$-valued bilinear form on $L_{K}\coloneqq L\otimes K$. Denote by $L^{\vee}\coloneqq\operatorname{Hom}_{R}(L,R)$ the dual of $L$. We have natural embeddings $L\hookrightarrow L^{\vee}\hookrightarrow L_{K}$ with the first map sending $x$ to $\phi(x,\cdot)$. The discriminant group of $L$ is by definition the quotient $A_{L}:=L^{\vee}/L$, which is a finite abelian group.
Definition  2.1.
Given two ($R$-)lattices $L_{1}$ and $L_{2}$, and an embedding $i:L_{1}\hookrightarrow L_{2}$
  • (1)
    $i$ is said to be primitive, if $L_{2}/L_{1}$ is a torsion-free $R$-module,
  • (2)
    $i$ is said to be a saturation of $L_{1}$, if $\operatorname{rank}L_{1}=\operatorname{rank}L_{2}$.
A lattice over $R=\mathbb{Z}$ is called an integral lattice; for a prime $p$, a lattice over $R=\mathbb{Z}_{p}$ is called a $p$-adic lattice. An integral (or $2$-adic) lattice $L$ is said to be even if for any $v\in L$, its norm $\phi(v,v)$ is even. Throughout this paper, we only encounter even lattices while over $\mathbb{Z}$ or $\mathbb{Z}_{2}$. An integral lattice $L$ is said to be unimodular if $A_{L}=0$.
2.2. Finite Quadratic Forms
We give a brief summary of finite quadratic forms following Nikulin [Nik79a, §$\S$ 1].
Definition  2.2.
  • (I)
    A finite symmteric bilinear form consists of a finite abelian group $A$ and a symmetric $\mathbb{Z}$-bilinear map $b:A\times A\rightarrow\mathbb{Q}/\mathbb{Z}$.
  • (II)
    A finite quadratic form consists of a finite abelian group $A$ and a function $q:A\rightarrow\mathbb{Q}/2\mathbb{Z}$ satisfying the following properties:
    1. (1)
      $q(na)=n^{2}q(a)$, for $n\in\mathbb{Z},a\in A$; and
    2. (2)
      if $b\colon A\times A\to\mathbb{Q}/\mathbb{Z}$ is defined by $b(a,b):=\frac{1}{2}(q(a+b)-q(a)-q(b))\in\mathbb{Q}/\mathbb{Z}$, then $(A,b)$ is a finite symmetric bilinear form.
    We call $b$ the bilinear form of $q$. We say $q$ is non-degenerate if $b$ is.
Let $(L,\phi)$ be an even integral lattice. We define $q_{L}$ to be a function on $A_{L}$ sending $[x]$ to $[\phi(x,x)]$ for $x\in L^{\vee}$. The pair $(A_{L},q_{L})$ is called the discriminant form associated with $L$. It is straightforward to check that $(A_{L},q_{L})$ is a non-degenerate finite quadratic form. We have the following decomposition:
(1) $\mathbb{Q}/\mathbb{Z}\cong\oplus_{p}\mathbb{Q}_{p}/\mathbb{Z}_{p},\mathbb{Q}/2\mathbb{Z}\cong(\oplus_{p\neq 2}\mathbb{Q}_{p}/\mathbb{Z}_{p})\oplus\mathbb{Q}_{2}/2\mathbb{Z}_{2}$
Proposition  2.3.
[Nik79a, §1, $2^{\circ}$] Let $(A,q)$ be a finite quadratic form. We decompose $A$ into a direct sum of its $p$-part
$A=\oplus_{p}A_{p}$
and $q_{p}=q|_{A_{p}}$. Then $(A,q)\cong\oplus_{p}(A_{p},q_{p})$. Via the decomposition (1), $q_{p}$ takes values in $\mathbb{Q}_{p}/\mathbb{Z}_{p}$ for $p\neq 2$, and $q_{2}$ takes values in $\mathbb{Q}_{2}/2\mathbb{Z}_{2}$.
For any finite abelian group $A$, we denote by $\ell(A)$ the minimal number of generators of $A$, and denote $\ell_{p}(A)=\ell(A_{p})$ for any prime $p$.
For the discriminant form $q_{L}$ of $L$, we define its signature $\tau(q_{L})$ as the signature of $L$ modulo $8$. By [Nik79a, Theorem 1.3.3], $\tau(q_{L})$ only depends on the finite quadratic form $q_{L}$.
2.3. Conway–Sloane Symbol
We take $R=\mathbb{Z}_{p}$.
Following Nikulin [Nik79a, §1, $8^{\circ}$], we denote the one-dimensional $\mathbb{Z}_{p}$-lattice $(\theta p^{k})$ by $K^{(p)}_{\theta,k}$, where $\theta\in\mathbb{Z}_{p}^{\times}$. And when $p=2$, we put:
$U^{(2)}_{k}=\begin{pmatrix}0&2^{k}\\ 2^{k}&0\end{pmatrix},V^{(2)}_{k}=\begin{pmatrix}2^{k+1}&2^{k}\\ 2^{k}&2^{k+1}\end{pmatrix}$
We denote the semigroup of isomorphism classes of p-adic lattices with respect to $\oplus$” by $\mathrm{Qu}(\mathbb{Z}_{p})$.
Proposition  2.4.
[Nik79a, Proposition 1.8.3]
  1. (1)
    if $p\neq 2$, $\mathrm{Qu}(\mathbb{Z}_{p})$ is generated by $K^{(p)}_{\theta,k}$;
  2. (2)
    if $p=2$, $\mathrm{Qu}(\mathbb{Z}_{2})$ is generated by $K^{(2)}_{\theta,k}$, $U^{(2)}_{k}$ and $V^{(2)}_{k}$.
Remark  2.5.
Given a $p$-adic lattice, its expression as a direct sum of $K^{(p)}_{\theta,k}$ (and $U^{(2)}_{k}$, $V^{(2)}_{k}$ for $p=2$) may not be unique, see [Nik79a, §1, $8^{\circ}$].
When $p$ is odd, let $(L,\phi)$ be a p-adic lattice. Then by Proposition 2.4, there exists a $\mathbb{Z}_{p}$-basis of $L$, such that with respect to this basis $\phi$ is a diagonal matrix of the following form:
$\operatorname{diag}(p^{s_{1}}M_{1},\cdots,p^{s_{t}}M_{t}).$
Here $0\leq s_{1}<s_{2}<\cdots<s_{t}$, each $M_{i}$ is a diagonal matrix of rank $r_{i}$, with diagonal terms in $\mathbb{Z}_{p}^{\times}$.
We define the Conway–Sloane symbol of $L$ as $\prod\limits_{s_{i}>0}(p^{s_{i}})^{\epsilon_{i}r_{i}}$ where
$\epsilon_{i}=\begin{cases}+,\text{if}\det(M_{i})\text{ is a square in }\mathbb{Z}_{p},\\ -,\text{otherwise}.\end{cases}$
For $p=2$, let $(L,\phi)$ be a $2$-adic even lattice. The Conway–Sloane symbol is more complicated. By Proposition 2.4 and [Nik79b, Proposition 1.8.2], there exists a $\mathbb{Z}_{2}$-basis of $L$, such that with respect to this basis $\phi$ is a block diagonal matrix of the following form:
$\operatorname{diag}(2^{s_{1}}M_{1},\cdots,2^{s_{t}}M_{t}).$
Here $M_{i}$ is either a diagonal matrix with diagonal terms in $\mathbb{Z}_{2}^{\times}$, or a block diagonal matrix with diagonal terms $U_{0}^{(2)}$ and $V_{0}^{(2)}$. In the former case we call $M_{i}$ of odd type, and for the latter case we call $M_{i}$ of even type. Suppose $M_{i}$ has rank $r_{i}$, which does not depend on the choices of basis. The Conway–Sloane symbol is defined as $\prod\limits_{s_{i}>0}{(2^{s_{i}})^{\epsilon_{i}r_{i}}_{t_{i}}}$ where
$\epsilon_{i}=\begin{cases}+,\text{if}\det(M_{i})\equiv\pm 1\;(\mathrm{mod}\;8),\\ -,\text{otherwise},\end{cases}$
and
$t_{i}=\begin{cases}\mathrm{Tr}(M_{i})\;(\mathrm{mod}\;8),\text{ if }M_{i}\text{ is of odd type},\\ \operatorname{II},\text{otherwise}.\end{cases}$
For an even integral lattice $L$, putting together the Conway–Sloane symbol of $L$ as a $p$-adic lattice for all possible $p$, we obtain the Conway–Sloane symbol of $L$.
If two even integral lattices $L$ and $L^{\prime}$ have same Conway–Sloane symbol, then the discriminant forms $(A_{L},q_{L})$ and $(A_{L^{\prime}},q_{L^{\prime}})$ are isomorphic. In other words, Conway–Sloane symbol determines the discriminant form up to isomorphism.
2.4. Niemeier Lattices and Leech Pairs
From now on, lattices are always integral if not stated otherwise.
Definition  2.6.
A Niemeier lattice is an even unimodular negative lattice with rank 24. There are 24 Niemeier lattices. Among the 24 Niemeier lattices, there is a unique one called the Leech lattice, denoted by $\mathbb{L}$, which is rootless (namely, $\mathbb{L}$ does not contain elements with norm -2).
The Leech lattice will play an important role in our study of finite symplectic automorphism groups of supersingular K3 surfaces.
Definition  2.7.
We denote the automorphism group of the Leech lattice $\mathbb{L}$ by $\operatorname{Co}_{0}$. And we put $\operatorname{Co}_{1}=\operatorname{Co}_{0}/\{\pm 1\}$. They are called Conway groups. The Conway group $\operatorname{Co}_{1}$ is a sporadic simple group.
Definition  2.8.
For a finite group $G$ acting faithfully on a negative definite even lattice $S$, the pair $(G,S)$ is called a Leech pair, if $S$ is rootless, the induced action of $G$ on $A_{S}$ is trivial, and the action has no non-zero fixed vectors.
The following criterion is crucial for us.
Proposition  2.9.
Given a Leech pair $(G,L)$ with $\operatorname{rank}(L)+\ell(A_{L})\leq 24$, then $(G,L)$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. In particular, $G$ is a subgroup of $\operatorname{Co}_{0}$.
This proposition is originally proved in [GHV12, appendix], see also [Huy16b, Proposition 2.2] and [LZ22, Proposition 3.4].
2.5. Some Lemmas on Lattice Embedding
We here collect some lemmas which will be frequently used in our later discussion.
Lemma  2.10.
Let $M\hookrightarrow N$ be a primitive embedding of two even lattices $N,M$, then $\operatorname{rank}(M)+\ell_{q}(A_{M})\leq\operatorname{rank}(N)+\ell_{q}(A_{N})$ for any prime number $q$. Moreover, $\operatorname{rank}(M)+\ell(A_{M})\leq\operatorname{rank}(N)+\ell(A_{N})$.
Proof.
Take a primitive embedding of $N$ into an even unimodular lattice $L$. Denote by $N_{L}^{\perp},M_{L}^{\perp}$ the orthogonal complements of $N,M$ in $L$ respectively. Then we have a primitive embedding $N^{\perp}_{L}\hookrightarrow M^{\perp}_{L}$. Let $K$ be the orthogonal complement of $N^{\perp}_{L}$ in $M^{\perp}_{L}$. Then we have a saturation $N^{\perp}_{L}\oplus K\hookrightarrow M^{\perp}_{L}$. This implies that:
$\ell_{q}(A_{M^{\perp}_{L}})\leq\ell_{q}(A_{N^{\perp}_{L}}\oplus A_{K}).$
Combining with $A_{M^{\perp}_{L}}\cong A_{M}$, $A_{N^{\perp}_{L}}\cong A_{N}$ and $\ell(A_{K})\leq\operatorname{rank}(K)=\operatorname{rank}(N)-\operatorname{rank}(M)$ we conclude $\operatorname{rank}(M)+\ell_{q}(A_{M})\leq\operatorname{rank}(N)+\ell_{q}(A_{N})$. Replace $l_{q}$ by $l$, the above argument still holds and we conclude $\operatorname{rank}(M)+\ell(A_{M})\leq\operatorname{rank}(N)+\ell(A_{N})$. ∎
Let $G$ be a group acting on a lattice $L$. Recall that $L^{G}$ is the $G$-invariant sublattice of $L$ and $S_{G}(L)$ is the orthogonal complement of $L^{G}$ in $L$, which is called the covariant lattice of the action.
Lemma  2.11.
Suppose a group $G$ acts on a lattice $L$, such that the induced action of $G$ on $A_{L}$ is trivial. Then $G$ acts trivially on $A_{S_{G}(L)}$.
Proof.
Since $S_{G}(L)\hookrightarrow L$ is a primitive embedding, we have a surjection $L^{\vee}\twoheadrightarrow(S_{G}(L))^{\vee}$. In fact, we have:
$S_{G}(L)\hookrightarrow L\hookrightarrow L^{\vee}\twoheadrightarrow(S_{G}(L))^{\vee}$
Given an element in $A_{S_{G}(L)}$, we can choose a representative $x\in L^{\vee}$. Since $G$ acts trivially on $A_{L}$, we have $gx-x\in L$ for any $g\in G$. Let $\phi$ be the bilinear form on $L$. For any $y\in L^{G}$, we have $\phi(gx-x,y)=0$, which implies that $gx-x\in S_{G}(L)$, i.e. $g[x]=[x]\in A_{S_{G}(L)}$. ∎
Lemma  2.12.
Suppose $M\hookrightarrow L$ is an embedding of even lattices. Suppose $G<\operatorname{Aut}(M)$ is a finite subgroup such that the induced action on $A_{M}$ is trivial. Then there exists an action of $G$ on $L$ extending the action of $G$ on $M$, such that any vector of $L$ orthogonal to $M$ is fixed by the action. Moreover, the induced action of $G$ on $A_{L}$ is trivial.
Proof.
We may assume $M$ is primitive in $L$ via replacing it by its saturation in $L$. Take a primitive embedding of $L$ into a unimodular even lattice $N$. The action of $G$ on $M$ extends to an action on $N$ that restricts to identity on the orthogonal complement of $M$ in $N$. Then the restriction of this action to $L$ gives the required action. ∎
2.6. Supersingular K3 Lattices
We begin with the following proposition, in which the lattice encodes the properties of the Néron-Severi lattice of a supersingular K3 surface.
Proposition-Definition  2.13.
[RS81, §1] For an odd prime $p$, and $1\leq\sigma\leq 10$, there exists a unique lattice, denoted as $N_{p,\sigma}$, with the following properties:
  • (1)
    The lattice $N_{p,\sigma}$ is even;
  • (2)
    The signature of $N_{p,\sigma}$ is (1,21);
  • (3)
    The discriminant group $A_{N_{p,\sigma}}$ is isomorphic to $(\mathbb{Z}/p\mathbb{Z})^{2\sigma}$.
And we call $N_{p,\sigma}$ the supersingular K3 lattice in characteristic $p$ and with Artin invariant $\sigma$.
In the following proposition, we recall the calculation of the Conway–Sloane symbol of a supersingular K3 lattice.
Proposition  2.14.
Let $N_{p,\sigma}$ be a supersingular K3 lattice as above. Then we have:
  1. (1)
[…]
    $q_{N_{p,\sigma}}=\begin{cases}p^{+2\sigma},\text{if}\;\sigma\;\text{is odd},\\ p^{-2\sigma},\text{if}\;\sigma\;\text{is even}.\end{cases}$
  2. (2)
[…]
Proof.
[…]
(See [Nik79a, Proposition 1.11.2] or [NT98, §2.1]). ∎
Corollary  2.15.
There is no maximal isotropic subspace of $(A_{N_{p,\sigma}},q_{N_{p,\sigma}})$ of dimension $\sigma$.
Proof.
We argue by contradiction. If such a maximal isotropic subspace exists, then we have:
$q_{N_{p,\sigma}}=\begin{cases}p^{+2\sigma},\text{if}\;(-1)^{\sigma}\;\text{is a quadratic residue},\\ p^{-2\sigma},\text{if}\;(-1)^{\sigma}\;\text{is not a quadratic residue}.\end{cases}$
Hence the corollary follows from the Proposition 2.14. ∎
3. Supersingular K3 Surfaces
In this section, we first collect some results on supersingular K3 surfaces. In §3.4, we specialize to superspecial K3 surfaces.
3.1. Néron–Severi Lattices
In this subsection, we focus on Néron–Severi lattices of supersingular K3 surfaces in odd characteristics. Let $X$ be a K3 surface in characteristic $p$, then we have:
$\operatorname{Pic}(X)\cong\operatorname{NS}(X)\cong\mathbb{Z}^{\oplus\rho(X)},1\leq\rho(X)\leq 22.$
It is known that $\operatorname{NS}(X)$ is an even lattice of signature $(1,\rho(X)-1)$, see [Huy16a, §17.2]. The K3 surface $X$ is called (Shioda-)supersingular if $\rho(X)=22$.
Theorem  3.1.
[Art74, RS81] Let $X$ be a supersingular K3 surface. Then there exists an integer $1\leq\sigma\leq 10$, the Artin invariant, such that:
$A_{\operatorname{NS}(X)}\cong(\mathbb{Z}/p\mathbb{Z})^{2\sigma}.$
By Proposition 2.13, for a supersingular K3 surface $X$ with Artin invariant $\sigma$, we have $\operatorname{NS}(X)\cong N_{p,\sigma}$.
Definition  3.2.
[Lie13, Definition 4.14] Let $\sigma\geq 1$ be an integer and $V$ be a $2\sigma$-dimensional $\mathbb{F}_{p}$-vector space endowed with a nondegenerate and nonneutral quadratic form. A subspace $K\subset V\otimes_{\mathbb{F}_{p}}k$ is called characteristic if
  • (a)
    $K$ is totally isotropic of dimension $\sigma$, i.e., all the vectors in $K$ are isotropic, and
  • (b)
    $K+(Id\otimes\operatorname{Fr}_{k})(K)$ is of dimension $\sigma+1$. Here $\operatorname{Fr}_{k}:k\rightarrow k$ is the Frobenius map.
3.2. Crystalline Torelli Theorem for Supersingular K3 Surfaces
We review the crystalline Torelli theorem for marked supersingular $K3$ surfaces.
Let $X$ be a supersingular K3 surface with Artin invariant $\sigma$. Next we write $N=N_{p,\sigma}$ for short. An $N$-marking of $X$ is a lattice isomorphism $\phi:N\rightarrow\operatorname{NS}(X)$. Let $(X,\phi),(Y,\psi)$ be two $N$-marked K3 surfaces. An isomorphism between $(X,\phi)$ and $(Y,\psi)$ is an isomorphism $\theta:X\rightarrow Y$ such that $\phi=\theta^{*}\circ\psi$. We define:
$\ker(X,\phi):=\ker(c_{k}:N\otimes k\rightarrow H^{2}_{dR}(X/k)).$
Here $c_{k}$ is the composition of $N\otimes k\rightarrow\operatorname{NS}(X)\otimes k$ and $\operatorname{NS}(X)\otimes k\rightarrow H^{2}_{crys}(X/W)\otimes k\cong H^{2}_{dR}(X/k)$. The second map is induced by the crystalline Chern class map. We put $K(X,\phi)\coloneqq(id\otimes\operatorname{Fr}_{k})^{-1}(\ker(X,\phi))$.
Proposition  3.3.
[Ogu79, Proposition 4.3][Ogu83, §1.8] Both $\ker(X,\phi)$ and $K(X,\phi)$ are characteristic subspaces of $(pN^{\vee}/pN)\otimes_{\mathbb{F}_{p}}k$.
We denote the set of roots in $N$ by $\Delta$. A line bundle $L$ on a K3 surface is ample if and only if $L\cdot L>0$, $L\cdot E>0$ for all effective divisors $E$ with $E\cdot E=-2$ and $h^{0}(L)\neq 0$. Denote by $\operatorname{Amp}(X)$ the ample cone of $X$. Then $\phi^{-1}(\operatorname{Amp}(X))$ is contained in a unique connected component of $\{x\in N\otimes\mathbb{R}|x\cdot x>0,x\cdot r\neq 0,\forall r\in\Delta\}$. For more details of the above statements, see [Ogu83, §1.10]. Now we can state the following version of the crystalline Torelli theorem using ample cones for $N$-marked supersingular K3 surfaces.
Theorem  3.4.
[Ogu83, BL18] Let $(X,\phi)$ and $(Y,\psi)$ be two $N$-marked supersingular K3 surfaces. If
$(K(X,\phi),\phi^{-1}(\operatorname{Amp}(X))=(K(Y,\psi),\psi^{-1}(\operatorname{Amp}(Y)),$
then there is a unique isomorphism $f:(X,\phi)\rightarrow(Y,\psi)$.
The theorem we state here appears as [Ogu83, Theorem II′′] for $p\geq 5$. It is equivalent to the crystalline Torelli theorem [Ogu83, Theorem II] which was proved there for $p\geq 5$ and by Bragg–Lieblich [BL18, Theorem 5.1.9] for $p\geq 3$. Thus Theorem 3.4 holds for all odd characteristics.
The following proposition is the key bridge for us to characterize finite groups of symplectic automorphisms of $X$.
Proposition  3.5.
For supersingular K3 surface $X$, we have the following isomorphism:
$\operatorname{Aut}^{s}(X)\cong\{g\in\mathrm{O}^{+}(\operatorname{NS}(X))|g\;\text{is identity on }A_{\operatorname{NS}(X)}\}$
here $\mathrm{O}^{+}$ means preserving the pairing and ample cone.
Proof.
Jang [Jan16, Proposition 3.1] shows that for any $g\in\operatorname{Aut}(X)$, $g$ acts on $A_{\operatorname{NS}(X)}\otimes_{\mathbb{F}_{p}}k$ trivially if and only if $g$ is symplectic. By Proposition 3.3, $K$ is contained in
$(p\operatorname{NS}(X)^{\vee}/p\operatorname{NS}(X))\otimes k\subset A_{\operatorname{NS}(X)}\otimes k.$
Then this proposition is a direct result of Theorem 3.4. ∎
3.3. A Lattice-Theoretic Interpretation of Symplectic Automorphism Groups
By [Nik79a, Theorem 1.14.4], there exists a unique primitive embedding:
$N_{p,\sigma}\hookrightarrow\operatorname{II}_{9,41}$
up to automorphisms of $\operatorname{II}_{9,41}$. Here the indices $9,41$ are simply chosen large enough to fulfill Nikulin’s criterion. We denote by $\mathbb{D}_{p,\sigma}$ the orthogonal complement of $N_{p,\sigma}$ in $\operatorname{II}_{9,41}$. Then $q_{\small{\mathbb{D}_{p,\sigma}}}=-q_{N_{p,\sigma}}$ and $\operatorname{sgn}(\mathbb{D}_{p,\sigma})=(8,20)$. Such a lattice $\mathbb{D}_{p,\sigma}$ is unique up to isomorphism and the orthogonal complement of any primitive embedding of $\mathbb{D}_{p,\sigma}$ in $\operatorname{II}_{9,41}$ is isomorphic to $N_{p,\sigma}$. We call $\mathbb{D}_{p,\sigma}$ the auxiliary lattice. The next proposition will be used to prove Theorem 3.8.
Proposition  3.6.
For a negative definite lattice $S$ with $\operatorname{rank}(S)\leq 21$, the following are equivalent:
  1. (1)
    there exists a primitive embedding $S\hookrightarrow N_{p,\sigma}$;
  2. (2)
    there exists an embedding $S\oplus\mathbb{D}_{p,\sigma}\hookrightarrow\operatorname{II}_{9,41}$ such that the image of $S$ is primitive.
Proof.
Assuming (1), then there is a primitive embedding $S\oplus\mathbb{D}_{p,\sigma}\hookrightarrow N_{p,\sigma}\oplus\mathbb{D}_{p,\sigma}$. By definition of $\mathbb{D}_{p,\sigma}$, we have a saturation $N_{p,\sigma}\oplus\mathbb{D}_{p,\sigma}\hookrightarrow\operatorname{II}_{9,41}$ with $N_{p,\sigma}$ primitive. The composition of the above two embeddings gives rise to an embedding $S\oplus\mathbb{D}_{p,\sigma}\hookrightarrow\operatorname{II}_{9,41}$ such that the image of $S$ is primitive.
Conversely, if (2) holds, we have $S\hookrightarrow\mathbb{D}_{p,\sigma}^{\perp}\cong N_{p,\sigma}$, which is apparently primitive. ∎
The following lemma is essential for further classifications:
Lemma  3.7.
Let $X$ be a supersingular K3 surface and $G\leq\operatorname{Aut}^{s}(X)$ a finite subgroup. Then $(G,S_{G}(X))$ is a Leech pair.
Proof.
Since $G$ is finite, there exists a $G$-invariant ample class on $X$. Since $\operatorname{sgn}(\operatorname{NS}(X))=(1,21)$, we know that the lattice $S_{G}(X)$ is negative definite. By Proposition 3.5, $G$ acts trivially on $A_{\operatorname{NS}(X)}$ , then $G$ acts trivially on $A_{S_{G}(X)}$ by Lemma 2.11. Since $S_{G}(X)$ is orthogonal to the ample class, it cannot contain any effective divisors. Therefore, there are no roots in $S_{G}(X)$. ∎
The following theorem is a criterion of Leech pairs arising from symplectic actions of finite groups on supersingular K3 surfaces:
Theorem  3.8.
Let $(G,S)$ be a Leech pair with $\operatorname{rank}(S)\leq 21$. The following are equivalent:
  1. (1)
    for each supersingular K3 surface $X/k$ with Artin invariant $\sigma$, there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
  2. (2)
    there exists a supersingular K3 surface $X/k$ with Artin invariant $\sigma$, such that there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
  3. (3)
    there exists a primitive embedding $S\hookrightarrow N_{p,\sigma}$.
Proof.
It is clear that (1) implies (2) and (2) implies (3). Next we show that (3) implies (1). We denote by $H$ the orthogonal complement of $S$ in $N_{p,\sigma}$. By Lemma 2.12, the action of $G$ on $S$ extends to $N_{p,\sigma}$ such that $G$ acts trivially on $A_{N_{p,\sigma}}$ and $H$.
Next, we show the existence of a vector $h\in H$ with positive norm, such that $h_{N_{p,\sigma}}^{\perp}$ contains no roots. Let $C$ be the set of elements in $H$ with positive norm. By [Ogu83, Proposition 1.10], the chambers in $(N_{p,\sigma})_{\mathbb{R}}$ with respect to all roots in $N_{p,\sigma}$ are open. Since $S$ contains no roots, there exists a chamber ${\mathcal{V}}$ such that ${\mathcal{V}}\cap C_{\mathbb{R}}\neq\{0\}$. Since ${\mathcal{V}}$ is open, we have ${\mathcal{V}}\cap C_{\mathbb{Q}}\neq\{0\}$. By taking an integral scalar of a nonzero element in ${\mathcal{V}}\cap H_{\mathbb{Q}}$, we obtain a primitive vector $h\in H$ with positive norm, such that $h_{N_{p,\sigma}}^{\perp}$ contains no root. By [Ogu83, Proposition 1.10], for each supersingular K3 surface $X/k$ with Artin invariant $\sigma$, there exists an isomorphism $f\colon N_{p,\sigma}\cong\operatorname{NS}(X)$, such that $f(h)\in\operatorname{NS}(X)$ is ample. The action on $N_{p,\sigma}$ naturally induces an action of $G$ on $\operatorname{NS}(X)$, such that $f(h)\in\operatorname{NS}(X)$ is fixed and the induced action on $A_{\operatorname{NS}(X)}$ is trivial. Therefore, the characteristic subspace is preserved, and by Proposition 3.5, we obtain an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, such that $(G,S_{G}(X))\cong(G,S)$. ∎
3.4. Superspecial K3 Surfaces
In this subsection, we discuss the case of superspecial K3 surfaces, i.e., $\sigma=1$. Maximal finite subgroups of $\operatorname{Aut}^{s}(X)$ for superspecial K3 surfaces over an algebraically closed field of any characteristic (including $2$) are obtained recently by Ohashi–Schütt [OS24]. For further use in this paper, we give a larger list (Table LABEL:table:_sublist_of_HM) consisting of all Höhn–Mason pairs $(G,S)$ which admit primitive embedding $S\hookrightarrow N_{p,1}$ for odd prime $p$. Here Höhn–Mason pair is defined as follows:
Definition  3.9.
A Höhn–Mason pair $(G,S)$ is a Leech pair with $G\leq\operatorname{Co}_{0}$ and $S=S_{G}(\mathbb{L})$, such that $G$ is maximal in the sense that $G$ contains all automorphisms of $S$ acting trivially on $A_{S}$, i.e., $G$ has no nontrivial enlargement.
Such kind of pairs have been classified by Höhn–Mason [HM16] and there are 290 of them. Now we relate finite groups of symplectic automorphisms of a superspecial K3 surface with Höhn–Mason pairs. Recall that for any Leech pair $(G,S)$, we let $\widetilde{G}$ be the group of all automorphisms of $S$ acting trivially on $A_{S}$. The enlargement $(\widetilde{G},S)$ is still a Leech pair.
Proposition  3.10.
Suppose $X$ is a superspecial K3 surface and $G\leq\operatorname{Aut}^{s}(X)$ is finite. Then the pair $(G,S_{G}(X))$ is a Leech pair and it can be enlarged to a Höhn–Mason pair $(\widetilde{G},S_{G}(X))$.
Proof.
By Lemma 3.7, the pair $(G,S_{G}(X))$ is a Leech pair. We have $\operatorname{rank}(\operatorname{NS}(X))+\ell(A_{\operatorname{NS}(X)})=24$. By Lemma 2.10 we have $\operatorname{rank}(S_{G}(X))+\ell(A_{S_{G}(X)})\leq 24$. By Proposition 2.9, $(G,S_{G}(X))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Thus $(\widetilde{G},S_{G}(X))$ is a Höhn–Mason pair. ∎
By Theorem 3.8, a Leech pair $(G,S)$ arises from a superspecial K3 surface in characteristic $p$ if and only if there exists a primitive embedding $S\hookrightarrow N_{p,1}$.
Theorem  3.11.
A Höhn–Mason pair $(G,S)$ admits primitive embedding $S\hookrightarrow N_{p,1}$ if and only if $(G,S,p)$ belongs to Table LABEL:table:_sublist_of_HM.
Proof.
Given a Höhn–Mason pair $(G,S)$, then $q_{S}$ is known. By Proposition 3.6, it suffices to check whether there exists an embedding $S\oplus\mathbb{D}_{p,\sigma}\hookrightarrow\operatorname{II}_{9,41}$ such that the image of $S$ is primitive. Suppose such an embedding exists.
[…]
Depending on $q_{S}$, there are at most three possibilities for $q_{\widetilde{S}}$, corresponding to $[\widetilde{S}:S\oplus\mathbb{D}_{p,1}]=1,p^{2},p^{4}$. For each possibility, we can use Nikulin’s existence theorem [Nik79a, Theorem 1.10.1] to check whether an even lattice of rank $(1,21-\operatorname{rank}(S))$ and with discriminant form $-q_{\widetilde{S}}$ exists. If not, then the corresponding possibility is excluded. If so, suppose $K$ is an even lattice of rank $(1,21-\operatorname{rank}(S))$ and $q_{K}=-q_{\widetilde{S}}$. Then there exists a saturation $\widetilde{S}\oplus K\hookrightarrow\operatorname{II}_{9,41}$. Since $S$ is primitive in $\widetilde{S}$, and $\widetilde{S}$ is primitive in $\operatorname{II}_{9,41}$, we obtain an embedding $S\oplus\mathbb{D}_{p,1}\hookrightarrow\operatorname{II}_{9,41}$ with image of $S$ primitive. ∎
Remark  3.12.
It is fine to replace the auxiliary lattice $\mathbb{D}_{p,1}$ by a negative even lattice of rank $4$ with same discriminant form.
Remark  3.13.
The results in Table LABEL:table:_sublist_of_HM with $\operatorname{rank}(S)=20,21$ have been obtained by Ohashi–Schütt [OS24], which is sufficient for their classification. The cases with $\operatorname{rank}(S)\leq 19$ will be useful to understand the size of finite groups of symplectic automorphisms of supersingular K3 surfaces with arbitrary Artin invariants, see Proposition 3.15 and §6. In the proof of Theorem 3.11, our description of the algorithm is brief. We refer to [OS24] for more details about the calculation.
Example  3.14.
Take $(G,q_{S})\cong(L_{3}(4).2,4_{3}^{-1}3^{-1}7^{-1})$ for an example (the corresponding number in the main list of Höhn–Mason is 170). In this case $p$ is supposed to be $3$ or $7$. Suppose $p=3$, then $\widetilde{S}$ is a saturation of $S\oplus\mathbb{D}_{3,1}$ of index $3$. We have $q_{S\oplus\mathbb{D}_{3,1}}=4_{3}^{-1}3^{-1}7^{-1}\oplus 3^{+2}=4_{3}^{-1}3^{-3}7^{-1}$, which implies that $q_{\widetilde{S}}=4_{3}^{-1}3^{+1}7^{-1}$, and $-q_{\widetilde{S}}=4_{5}^{-1}3^{-1}7^{+1}$. If $p=7$, we have $q_{S\oplus\mathbb{D}_{7,1}}=4_{3}^{-1}3^{-1}7^{-3}$ and $-q_{\widetilde{S}}=4_{5}^{-1}3^{+1}7^{-1}$. Notice that the rank one lattice $\langle 84\rangle$ has discriminant form $4_{5}^{-1}3^{+1}7^{-1}$. Therefore, $(G,q_{S})\cong(L_{3}(4).2,4_{3}^{-1}3^{-1}7^{-1})$ is realized for $p=7$ and not possible for $p=3$.
3.5. Artin Invariants $\sigma\geq 2$
In this subsection, let us take a glimpse at supersingular K3 surfaces with Artin invariants $\sigma\geq 2$. Let $X$ be such a supersingular K3. Then $\operatorname{NS}(X)\cong N_{p,\sigma}$.
Suppose $G\leq\operatorname{Aut}^{s}(X)$ is a finite subgroup. Then there is an induced action of $G$ on $N_{p,\sigma}$ with covariant lattice $S_{G}(N_{p,\sigma})$. Since the action of $G$ on $A_{N_{p,\sigma}}$ is trivial, there exists a saturation $N_{p,\sigma}\hookrightarrow N_{p,1}$ of index $p^{2\sigma-2}$, such that the action of $G$ on $N_{p,\sigma}$ extends to an action on $N_{p,1}$.
The covariant lattice $S_{G}(N_{p,1})$ is a saturation of $S_{G}(N_{p,\sigma})$. Both $S_{G}(N_{p,\sigma})$ and $S_{G}(N_{p,1})$ are negative definite. The lattice $S_{G}(N_{p,\sigma})$ has no roots, while $S_{G}(N_{p,1})$ may have. Denote by $R_{G}$ the root sublattice of $S_{G}(N_{p,1})$. Denote by $R_{G}^{\perp}$ the orthogonal complement of $R_{G}$ in $S_{G}(N_{p,1})$. Clearly, the action of $G$ on $S_{G}(N_{p,1})$ preserves $R_{G}$ and $R_{G}^{\perp}$. Thus we have a group homomorphism $G\to\operatorname{Aut}(R_{G})$. Let $G_{L}$ and $G_{R}$ be the kernel and image of this homomorphism, i.e., we have an exact sequence of groups:
(2) $1\longrightarrow G_{L}\longrightarrow G\longrightarrow G_{R}\longrightarrow 1$
The lattice $R_{G}^{\perp}$ contains no roots. The covariant sublattice $S_{G_{L}}(N_{p,1})$ is a primitive sublattice of $R_{G}^{\perp}$. Thus $(G_{L},S_{G_{L}}(N_{p,1}))$ is a Leech pair. Since $\operatorname{rank}(N_{p,1})+\ell(A_{N_{p,1}})=24$ and $S_{G_{L}}(N_{p,1})$ is a primitive sublattice of $N_{p,1}$, we have $\ell(A_{S_{G_{L}}(N_{p,1})})+\operatorname{rank}(S_{G_{L}}(N_{p,1}))\leq 24$ by Lemma 2.10. Then by Proposition 2.9, the Leech pair $(G_{L},S_{G_{L}}(N_{p,1}))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Moreover we have the following characterization of $(G_{L},S_{G_{L}}(N_{p,1}))$.
Proposition  3.15.
The enlargement $(\widetilde{G}_{L},S_{G_{L}}(N_{p,1}))$ of $(G_{L},S_{G_{L}}(N_{p,1}))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
Proof.
This is because $S_{G_{L}}(N_{p,1})$ is a primitive sublattice of $N_{p,1}$. ∎
Lemma  3.16.
The root lattice $R_{G}$ satisfies the following statements:
  1. (1)
    $(R_{G})^{G_{R}}=\{0\}$,
  2. (2)
    the sublattice generated by $pR_{G}$ and elements of the form $g\alpha-\alpha$ for $g\in G,\alpha\in R_{G}$ contains no roots.
Proof.
The first statement follows from that $R_{G}\subset S_{G}(N_{p,1})$.
For the second statement, we show that the sublattice is contained in $S_{G}(N_{p,\sigma})$. Since $N_{p,\sigma}\supset pN_{p,1}$, we have $S_{G}(N_{p,\sigma})\supset pS_{G}(N_{p,1})$. Thus $R_{G}\cap S_{G}(N_{p,\sigma})\supset pR_{G}$.
Since $G$ acts trivially on $A_{S_{G}(N_{p,\sigma})}$, the equivalence class of $g\alpha-\alpha$ in $A_{S_{G}(N_{p,\sigma})}$ is trivial for all $g\in G$ and $\alpha\in S_{G}(N_{p,1})$. This implies that $g\alpha-\alpha\in S_{G}(N_{p,\sigma})$. We conclude that the lattice generated by $pR_{G}$ and elements of the form $g\alpha-\alpha$ for $g\in G,\alpha\in R_{G}$ is contained in $S_{G}(N_{p,\sigma})$. Since there is no root in $S_{G}(N_{p,\sigma})$, we are done. ∎
Remark  3.17.
The automorphisms of root sytems of type $A,D,E$ will come into play. Later in §5, we will systematically study root systems which admit a finite group action satisfying the conditions in Lemma 3.16.
4. The Tame Case
In this section we give the proof of Theorem 1.1.
4.1. Proof of Theorem 1.1 for Supersingular K3 Surfaces
The next lemma will be used in the proof of Theorem 1.1 for supersingular K3 surfaces.
Lemma  4.1.
If $\gcd(|G|,p)=1$, then $\ell_{p}(A_{S_{G}(X)})=0$, and $S_{G}(X)\cong S_{G}(N_{p,1})$.
Proof.
Take any $[x]\in A_{S_{G}(X)}$ with $x\in S_{G}(X)_{\mathbb{Q}}$. The sum $\sum\limits_{g\in G}gx\in S_{G}(X)_{\mathbb{Q}}$ is fixed by $G$, hence equals to zero. On the other hand, since $G$ induces a trivial action on $A_{S_{G}(X)}$, we must have $gx-x\in S_{G}(X)$ for any $g\in G$. Therefore, $\sum\limits_{g\in G}(gx-x)\in S_{G}(X)$, which implies that $|G|x\in S_{G}(X)$ where $\gcd(|G|,p)=1$. We conclude that $\ell_{p}(A_{S_{G}(X)})=0$. Since $S_{G}(X)\subset S_{G}(N_{p,1})$ is of index a $p$-power, and $\gcd(|A_{S_{G}(X)}|,p)=1$, we must have $S_{G}(X)\cong S_{G}(N_{p,1})$. ∎
By Lemma 3.7, the pair $(G,S_{G}(X))$ is a Leech pair. We have $\operatorname{rank}(N_{p,1})+\ell_{q}(A_{N_{p,1}})=22$ if $q\neq p$. By Lemma 2.10 we have $\operatorname{rank}(S_{G}(N_{p,1}))+\ell_{q}(A_{S_{G}(N_{p,1})})\leq 22$. It is clear that $\operatorname{rank}(S_{G}(N_{p,1}))+\ell_{p}(A_{S_{G}(N_{p,1})})\leq 22$ since $\ell_{p}(A_{S_{G}(N_{p,1})})=0$. Therefore, $\operatorname{rank}(S_{G}(N_{p,1}))+\ell(A_{S_{G}(N_{p,1})})\leq 22$. From Lemma 4.1 and Proposition 2.9, the Leech pair $(G,S_{G}(X))$ is isomorphic to a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Since $S_{G}(X)$ admits a primitive embedding into $N_{p,1}$, the enlargement of $(G,S_{G}(X))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
We have a saturation $S_{G}(X)\oplus\operatorname{NS}(X)^{G}\subset\operatorname{NS}(X)$. By Lemma 4.1, $\ell_{p}(A_{S_{G}(X)})=0$, hence $\ell_{p}(\operatorname{NS}(X)^{G})=\ell_{p}(\operatorname{NS}(X))=2\sigma$. This implies that $\operatorname{rank}(\operatorname{NS}(X)^{G})\geq 2\sigma$. We conclude that $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$. Now the proof of Theorem 1.1 in the supersingular case is complete.
4.2. Proof of Theorem 1.1 for K3 Surfaces of Finite Height
In this subsection, we prove Theorem 1.1 for K3 surfaces of finite height $h$. We mainly use Jang [Jan17] to turn this into a problem over characteristic 0. For more on K3 surfaces of finite height, see [AM77] or [Lie16, §6] and references therein.
The following lemma is essentially due to Jang ([Jan17, Theorem 3.2, Corollary 3.4]):
Lemma  4.2.
Suppose $G\leq\operatorname{Aut}^{s}(X)$ is a finite subgroup such that $\gcd(|G|,p)=1$. Then we have the following.
  1. (1)
    The induced action of $G$ on $\operatorname{NS}(X)$ is faithful.
  2. (2)
    The induced action of $G$ on $A_{\operatorname{NS}(X)}$ is trivial.
  3. (3)
    $\operatorname{NS}(X)$ admits a primitive embedding into the K3 lattice $U^{3}\oplus E_{8}^{2}$, in particular we have $\operatorname{rank}(\operatorname{NS}(X))+\ell(A_{\operatorname{NS}(X)})\leq 22$.
Proof.
For any $g\in G$, by the proof of [Jan17, Corollary 3.4], there exists a Néron–Severi group preserving lifting $\mathscr{X}/W(k)$ of $X$, i.e. $\operatorname{NS}(\mathscr{X}_{\overline{K}})\cong\operatorname{NS}(X)$ and the action of $g$ on $X$ can be lifted to an action on $\mathscr{X}/W(k)$ compatible with the specialization map $\operatorname{NS}(\mathscr{X}_{\overline{K}})\cong\operatorname{NS}(X)$.
By Lieblich–Maulik [LM18, Theorem 2.1, Corollary 2.4] (see also Jang [Jan17, Page 1790]), we have an inclusion $\operatorname{Aut}(\mathscr{X}_{\overline{K}})\hookrightarrow\operatorname{Aut}(X)$. In particular, $\operatorname{Aut}(\mathscr{X})\hookrightarrow\operatorname{Aut}(X)$ where $\operatorname{Aut}(\mathscr{X})$ is the automorphism group of $\mathscr{X}$ as a $W(k)$-scheme.
Since $g$ is of finite order, then we know that its lifting, still denoted by $g$, is also of finite order (coprime with $p$). We denote the natural representation $\rho:\operatorname{Aut}(\mathscr{X})\rightarrow\operatorname{GL}(H^{0}(\mathscr{X},\omega_{\mathscr{X}}))=W(k)^{\times}$, then $\rho(g)$ acts as a root of unit in $W(k)^{\times}$ since $g$ of finite order. But $g$ is symplectic on the special fiber $X$, hence the eigenvalue is 1 in the residue field $k$. Then it has to be 1 in $W(k)$ since the order of $g$ is coprime with $p$. Therefore, $\rho(g)=Id$. Thus $g$ is also symplectic when restricted to $\mathscr{X}_{\overline{K}}$.
Next we show that $g$ acts trivially on $A_{\operatorname{NS}(\mathscr{X}_{\overline{K}})}$. Since such a conclusion is true over $\mathbb{C}$, we can use the ”Lefschetz principle” here to conclude over an algebraically closed field of characteristic 0. To be more precise, there exists a finitely generated field $E$ over $\mathbb{Q}$ contained in $\bar{K}$ and a K3 surfaces $X^{\prime}$ defined over $E$ with $X^{\prime}\times_{E}\overline{K}\cong\mathscr{X}_{\overline{K}}$ and $g$ acts on $X^{\prime}$ symplectically and of finite order. Since $E$ is finitely generated over $\mathbb{Q}$, there is also a field embedding $E\hookrightarrow\mathbb{C}$. We denote by $X^{\prime}_{\mathbb{C}}=X^{\prime}\times_{E}\mathbb{C}$. By [Huy16a, §17, Lemma 2.2], we have $\operatorname{NS}(\mathscr{X}_{\overline{K}})\cong\operatorname{NS}(X^{\prime}_{\overline{E}})\cong\operatorname{NS}(X^{\prime}_{\mathbb{C}})$, and the action of $g$ on $X^{\prime}_{\mathbb{C}}$ is symplectic. We then know that the induced action of $g$ on $H^{2}(X^{\prime}_{\mathbb{C}},\mathbb{Z})$ fixes all elements in the transcendental lattice $T(X^{\prime}_{\mathbb{C}})$. Therefore, the induced action of $g$ on $A_{\operatorname{NS}(X)}\cong A_{\operatorname{NS}(X^{\prime}_{\mathbb{C}})}\cong A_{T(X^{\prime}_{\mathbb{C}})}$ is trivial.
Each $g\in G$ acts faithfully on $\operatorname{NS}(X^{\prime}_{\mathbb{C}})$, hence also faithfully on $\operatorname{NS}(X)$. From the isomorphism $\operatorname{NS}(X)\cong\operatorname{NS}(X^{\prime}_{\mathbb{C}})$ we also know that $\operatorname{NS}(X)$ admits a primitive embedding into the K3 lattice. ∎
Applying Lemma 2.10 for the primitive sublattice $S_{G}(X)\subset\operatorname{NS}(X)$, we have
$\operatorname{rank}(S_{G}(X))+\ell(A_{S_{G}(X)})\leq 22.$
By Lemma 2.11, we know that the induced action of $G$ on $A_{S_{G}(X)}$ is trivial. Then it follows from Proposition 2.9 that $(G,S_{G}(X))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. It is clear from Lemma 4.2 that $S_{G}(X)$ also admits a primitive embedding into the K3 lattice.
Notice that $\operatorname{rank}(\operatorname{NS}(X))\leq 22-2h$ (see [AM77, Page 130]). Since $G$ is finite, we have $\operatorname{rank}(S_{G}(X))\leq 21-2h$.
Finally, since $\gcd(|G|,p)=1$, we know $\gcd(|A_{S_{G}(X)}|,p)=1$, similar as Lemma 4.1. We hence finish the proof of Theorem 1.1 for K3 surfaces of finite height.
5. $p$-Root Pairs
We start this section with the following definition motivated by Lemma 3.16.
Definition  5.1.
Let $H$ be a finite group acting faithfully on a root lattice $R$. We call $(H,R)$ a pseudo $p$-root pair if the lattice $R_{H}^{\#}$ generated by $pR$ and $\{g\alpha-\alpha\big{|}g\in H,\alpha\in R\}$ contains no roots. We call $(H,R)$ a $p$-root pair if it is a pseudo $p$-root pair, and in addition $R^{H}=\{0\}$.
We say a (pseudo) $p$-root pair $(H,R)$ is irreducible if $R$ is an irreducible root lattice (namely, the associated Dynkin diagram is connected). We will simply write $R^{\#}_{H}$ as $R^{\#}$ if no ambiguity caused.
In the following, we will first discuss irreducible $p$-root pairs and then general $p$-root pairs. In particular, in the last subsection, we prove the second statement in Theorem 1.2.
5.1. $p$-Root Pairs of Type $A$
We first study (pseudo) $p$-root pairs $(H,R)$ of type $A$. This means that $R$ is a root lattice of type $A_{m}$ for certain positive integer $m$. Then we can realize $R$ as the subgroup of the standard Euclidean space $\mathbb{R}^{m+1}$ generated by $v_{i}-v_{j}$. Here $(v_{1},\cdots,v_{m+1})$ is the standard basis of $\mathbb{R}^{m+1}$. The Weyl group $W(R)$ of $R=A_{m}$ can be identified with the permutation group $S_{m+1}$ of vectors $v_{1},\cdots,v_{m+1}$. The roots $\alpha_{i}=v_{i}-v_{i+1}$ for $1\leq i\leq m$ form a base of the root lattice.
Lemma  5.2.
Let $(H,R)$ be a pseudo $p$-root pair, with $R=A_{m}$. Then $H\cap W(A_{m})$ is a $p$-group, and for any nontrivial $g\in H\cap W(A_{m})$, it has no fixed element in $\{v_{1},\cdots,v_{m+1}\}$. In particular, it implies that if $H\cap W(A_{m})$ is nontrivial, then $p|(m+1)$.
Proof.
We first prove that $H\cap W(A_{m})$ is a $p$-group. Suppose there is an element $g\in H\cap W(A_{m})$ of order $q$ where $q$ is a prime and $q\neq p$. Without loss of generality, we assume that $g$ has a cycle $(v_{1},v_{2},\cdots,v_{q})$. Then
$\sum_{i=1}^{q}(g^{i}(\alpha_{1})-\alpha_{1})=-q\alpha_{1}\in R^{\#}.$
Since $p\alpha_{1}\in R^{\#},p\neq q$, we deduce $\alpha_{1}\in R^{\#}$, a contradiction. So $H\cap W(A_{m})$ is a $p$-group. Therefore, $\left|H\right|=p^{n}$ or $\left|H\right|=2\cdot p^{n}$ for certain integer $n\geq 0$.
Next, we prove that any nontrivial $g\in H\cap W(A_{m})$ has no fixed element in $\{v_{1},\cdots,v_{m+1}\}$. Otherwise we may choose an invariant element which we may denote as $v_{m+1}$. Since $g$ is nontrivial, there exists $v_{i}$, such that $g(v_{i}-v_{m+1})-(v_{i}-v_{m+1})$ is a root, a contradiction. The last statement hence follows. ∎
Corollary  5.3.
Let $(H,R)$ be a pseudo $p$-root pair, with $R=A_{m}$. For any $g\in H\cap W(A_{m})$, if $g$ is nontrivial, then all the orbits of $g$ have same length.
Proof.
In the proof of the above lemma, we have shown that for any element $g\in H\cap W(A_{m})$, $g$ has no fixed element in $\{v_{1},\cdots,v_{m+1}\}$. Apply this result to all the powers of $g$, then we can conclude. ∎
Remark  5.4.
This is a quite restrictive condition on $H$. However, $H$ needs not to be cyclic. For example,
$H=\{id,(12)(34),(13)(24),(14)(23)\}\subset S_{4}.$
Cycles of any nontrivial element have the same length, but $H$ is not cyclic.
For irreducible $p$-root pairs of type A, we have a stronger conclusion:
Proposition  5.5.
Given a $p$-root pair $(H,R)$ with $R=A_{m}$, then $(m+1)\big{|}|H\cap W(A_{m})|$. In particular, $m+1$ is a power of $p$.
Proof.
If $H\leq W(A_{m})$, we claim that $H$ acts transitively on $\{v_{1},\cdots,v_{m+1}\}$. Otherwise, there exist two orbits $U,V\subset\{v_{1},\cdots,v_{m+1}\}$ for the action of $H$. Take $v_{i}\in U,v_{j}\in V$. Consider the root $v_{i}-v_{j}$. Since $R^{H}=\{0\}$, we have $\sum\limits_{g\in H}g(v_{i}-v_{j})=0$. Then
$\sum\limits_{g\in H}gv_{i}=\sum\limits_{g\in H}gv_{j},$
But this is impossible since the left side is a positive combination of elements in $U$, while the left is a positive combination of elements in $V$. Therefore, $G$ acts transitively on $\{1,2,\cdots,m+1\}$ and hence $(m+1)\big{|}|H\cap W(A_{m})|$.
If $H\not\subset W(A_{m})$. Let $H^{\prime}=H\cap W(A_{m})$. Then $|H|=2|H^{\prime}|$ is even. So there exists an element of order $2$ in $H$, which is not in $W(A_{m})$. So we can choose an involution $w\in W(A_{m})$ such that $-w\in H$. Then $H=H^{\prime}\cup(-wH^{\prime})$.
Since $R^{H}=\{0\}$, then for $x\in R$, we have $\sum\limits_{g\in H}g\cdot x=0$. We claim that $H^{\prime}$ acts transitively on $\{1,\cdots,m+1\}$. Otherwise, we consider two orbits of $H^{\prime}$, say $U,V$ and take $v_{i}\in U,v_{j}\in V$. Let $x=v_{i}-v_{j}$, then we have:
$\sum_{h\in H^{\prime}}h\cdot v_{i}-\sum_{h\in H^{\prime}}wh\cdot v_{i}-\sum_{h\in H^{\prime}}h\cdot v_{j}+\sum_{h\in H^{\prime}}wh\cdot v_{j}=0$
Since $H^{\prime}$ is normal in $H$, it is straightforward to see that $w$ sends an $H^{\prime}$-orbit to an $H^{\prime}$-orbit. By the equality above, we know that $w(U)=U,w(V)=V$. But $H^{\prime}$ is a $p$-group, thus both $\left|U\right|,\left|V\right|$ are odd, and since $w$ is of order 2, there exists $v_{i^{\prime}}\in U,v_{j^{\prime}}\in V$ such that $w(v_{i^{\prime}})=v_{i^{\prime}},w(v_{j^{\prime}})=v_{j^{\prime}}$. Then
$-w(v_{i^{\prime}}-v_{j^{\prime}})-(v_{i^{\prime}}-v_{j^{\prime}})=-2(v_{i^{\prime}}-v_{j^{\prime}})\in R^{\#}.$
Since $p(v_{i^{\prime}}-v_{j^{\prime}})\in R^{\#}$, we have $v_{i^{\prime}}-v_{j^{\prime}}\in R^{\#}$, a contradiction. Thus we know $H^{\prime}$ acts transitively on $\{1,2,\cdots,m+1\}$, which implies $(m+1)\mid\left|H\cap W(A_{m})\right|$. ∎
Proposition  5.6.
If $(H,R)$ is a $p$-root pair with $R=A_{m}$, then $(H\cap W(A_{m}),R)$ is also a $p$-root pair.
Proof.
We put $H^{\prime}=H\cap W(A_{m})$ and we only need to show that $R^{H^{\prime}}=\{0\}$. From the proof of Proposition 5.5, we know that $H^{\prime}$ acts transitively on $\{1,2,\ldots,m+1\}$. For $x\in R^{H^{\prime}}$, we write $x=\sum\limits_{i=1}^{m+1}x_{i}v_{i}$. By the transitivity of $H^{\prime}$-action, and $x\in R^{H^{\prime}}$, we know that $x_{i}$ are the same for $11\leq i\leq m+$ 1, hence they are all 0. ∎
Remark  5.7.
This may not hold in general. See Lemma 5.14.
5.2. $p$-Root Pairs of Type $D_{m}$ with $m\geq 5$
In this subsection we study (pseudo) $p$-root pairs of type $D$. Let $v_{1},\cdots,v_{m}$ be the standard basis for the standard Euclidean space $\mathbb{R}^{m}$. Recall that the root lattice $D_{m}$ ($m\geq 4$) can be realized as the subgroup of $\mathbb{R}^{m}$ generated by ±$\pm v_{i}\pm v_{j}$, $1\leq i<j\leq m$. We choose the set of simple roots as $\{\alpha_{1},\cdots,\alpha_{m}\}$ with $\alpha_{i}=v_{i}-v_{i+1},1\leq i\leq m-1,\alpha_{m}=v_{m-1}+v_{m}$
For $g\in S_{m}$ acting on $\{v_{1},\cdots,v_{m}\}$, and $\epsilon_{1},\cdots,\epsilon_{m}\in\{\pm 1\}$, there is an element in $\operatorname{GL}(\mathbb{R},m)$ given by $v_{i}\mapsto\epsilon_{i}g(v_{i})$. These elements form a finite subgroup $S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m}$ of $\operatorname{Aut}(D_{m})$. When $m\geq 5$, we have an equality $\operatorname{Aut}(D_{m})=S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m}$. When $m=4$, we have $[\operatorname{Aut}(D_{4}):S_{4}\ltimes(\mathbb{Z}/2\mathbb{Z})^{4}]=3$.
The Weyl group $W(D_{m})$ of $D_{m}$ is isomorphic to $S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m-1}$. For any element $g\in S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m}$, we denote by $\overline{g}$ its part in $S_{m}$.
Lemma  5.8.
Suppose $(H,R)$ is a pseudo $p$-root pair of type $D_{m}$ for $m\geq 4$. Then for any $g\in H\cap(S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m})$, we have $\overline{g}=id$. And there exists at most one $i\in\{1,\cdots,m\}$ such that $g(v_{i})=-v_{i}$.
Proof.
If there is an element $g\in H\cap(S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m})$ whose image $\overline{g}$ in $S_{m}$ has an orbit of length at least 3, then without loss of generality, we assume that $\overline{g}(v_{1})=v_{2}$ and $\overline{g}(v_{2})=v_{3}$.
If $g(v_{1})=v_{2}$, then $g(v_{1}+v_{2})-(v_{1}+v_{2})$ is a root. If $g(v_{1})=-v_{2}$, then $g(v_{1}-v_{2})-(v_{1}-v_{2})$ is a root. Hence we rule out this possibility.
If $g\in H\cap(S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m})$ such that $\overline{g}$ has a cycle of length two. We assume $\overline{g}(v_{1})=v_{2}$ and $\overline{g}(v_{2})=v_{1}$.
If $g(v_{1})=v_{2}$ and $g(v_{2})=v_{1}$, we have $g(v_{1}-v_{2})-(v_{1}-v_{2})=-2(v_{1}-v_{2})$. Since $p(v_{1}-v_{2})\in R^{\#}$, we deduce $v_{1}-v_{2}\in R^{\#}$, a contradiction.
If $g(v_{1})=v_{2}$ and $g(v_{2})=-v_{1}$, then $g(v_{1}-v_{2})-(v_{1}-v_{2})=2v_{2}\in R^{\#}$, and $g(v_{1}+v_{2})-(v_{1}+v_{2})=-2v_{1}\in R^{\#}$, thus $2(v_{1}-v_{2})\in R^{\#}$, still a contradiction. The remaining cases are similar. We conclude that $\overline{g}=id$.
Suppose $g$ has more than one sign changes. Suppose $g(v_{1})=-v_{1},g(v_{2})=-v_{2}$. Then $g(v_{1}-v_{2})-(v_{1}-v_{2})=-2(v_{1}-v_{2})\in R^{\#}$. Since $p(v_{1}-v_{2})\in R^{\#}$, we have $v_{1}-v_{2}\in R^{\#}$, a contradiction. The lemma follows. ∎
Proposition  5.9.
Let $(H,R)$ be a pseudo $p$-root pair with $R=D_{m}$ for $m\geq 4$, then $H\cap W(R)=\{id\}$.
Proof.
We have $W(D_{m})=S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m-1}<S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m}$ consisting of elements with even number of sign changes. Suppose $g\in H\cap W(A_{m})$, then by Lemma 5.8, we have $\overline{g}=id$ and $g$ has at most one sign change. The only possibility is $g=id$. The proposition follows. ∎
Proposition  5.10.
Suppose $R=D_{m},m\geq 5$, then $(H,R)$ is a pseudo $p$-root pair if and only if:
  1. (1)
    $H=\{id\}$ or,
  2. (2)
    $H=\langle g\rangle$ with $g(v_{i})=-v_{i}$ for a unique $i$, and $g(v_{j})=v_{j}$ for $j\in\{1,\cdots,m\}-\{i\}$.
Moreover, there is no $p$-root pair of type $D_{m}$ with $m\geq 5$.
Proof.
If $m\geq 5$, we have $\operatorname{Aut}(D_{m})=S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m}$. By Lemma 5.8, any element $g\in H$ belongs to the two possibilities. It is straightforward to verify that the listed $H$ indeed form pseudo $p$-root pairs.
The last statement follows if we add one more restriction that $R^{H}=\{0\}$. ∎
Proposition  5.11.
Any $p$-root pair $(H,R)$ does not have a component of type $D_{m}$ with $m\geq 5$.
Proof.
This directly follows from Lemma 5.22 and Proposition 5.10. ∎
5.3. $p$-Root Pair of Type $D_{4}$
We have that $\operatorname{Aut}(D_{4})=S_{3}\ltimes W(D_{4})$. Consider the outer automorphisms: $x:\alpha_{3}\mapsto\alpha_{4}\mapsto\alpha_{1}\mapsto\alpha_{3},\alpha_{2}\mapsto\alpha_{2}$, $y:v_{4}\rightarrow-v_{4}$, i.e., $y:\alpha_{3}\mapsto\alpha_{4}$. Then the outer automorphism group $S_{3}$ is generated by $x,y$.
We take an element $g\in\operatorname{Aut}(D_{4})$ such that $g\alpha_{2}=-2\alpha_{2}-\alpha_{1}-\alpha_{3}-\alpha_{4},g(\alpha_{i})=\alpha_{i}+\alpha_{2},i=1,3,4$. Spell this out, we have $g:-v_{1}\mapsto v_{3}\mapsto v_{2}$. In other words $g=(v_{1}\mapsto-v_{1})\circ(132)\circ(v_{1}\mapsto-v_{1})$.
Lemma  5.12.
$\langle g\rangle\times\langle x\rangle$ is a Sylow 3-subgroup of $\operatorname{Aut}(D_{4})$. And any subgroup of $\operatorname{Aut}(D_{4})$ of order 3 is conjugate to one of the follows: $\langle g\rangle,\langle x\rangle,\langle gx\rangle$.
Proof.
We first verify that $gx=xg,g^{3}=id$, and $\langle g,x\rangle\cong\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$ is a Sylow 3-subgroup of $\operatorname{Aut}(D_{4})$. Any subgroup of $\operatorname{Aut}(D_{4})$ of order 3 has to be conjugate to one of $\langle g\rangle,\langle x\rangle,\langle gx\rangle,\langle gx^{2}\rangle$.
Notice that $g$ acts identically on $\langle v_{4},v_{1}-v_{2}-v_{3}\rangle$ which contains no roots, while $x$ acts identically on $\{\alpha_{2},\alpha_{1}+\alpha_{3}+\alpha_{4}\}$. We can calculate that $gx,gx^{2}$ do not fix any elements in $R$. Hence $\langle x\rangle,\langle g\rangle,\langle gx\rangle$ are not conjugate to each other in $\operatorname{Aut}(D_{4})$. Recall that $y\in\operatorname{Aut}(D_{4}):v_{4}\mapsto-v_{4}$, we can verify that $y(gx)y^{-1}=gx^{2}$. ∎
Proposition  5.13.
If $p\geq 5$, there is no $p$-root pair of type $D_{4}$.
Proof.
By Proposition 5.9, we know that for a pseudo $p$-root pair $(H,D_{4})$, $H$ is isomorphic to a subgroup of $\langle x,y\rangle=S_{3}$ induced by the following exact sequence:
$1\rightarrow W(D_{4})\rightarrow\operatorname{Aut}(D_{4})\rightarrow S_{3}\rightarrow 1$
We can verify that $R^{\#}$ has root $\alpha_{2}$ for $(\langle g\rangle,D_{4})$. Hence $(\langle g\rangle,D_{4})$ is not a pseudo-root pair. For $(\langle gx\rangle,D_{4})$,
$R^{\#}=\langle 3D_{4},2\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{1}+\alpha_{3}+\alpha_{4},pD_{4}\rangle$
Hence if $p\neq 3$, $R^{\#}$ has roots for $(\langle gx\rangle,D_{4})$.
We first assume that $|H|=3$, then by the previous arguments, since $p\geq 5$, $H$ has to be conjugate to $\langle x\rangle$ and then $R^{H}\neq 0$. Now suppose that $|H|=6,\langle x\rangle\subset H$ is a normal subgroup, and let $\gamma\in H$ be an order 2 element. Since $\langle x\rangle$ is normal, then $\gamma$ acts on $R^{x}$ which contains $\mathbb{Z}\alpha_{2}+\mathbb{Z}(\alpha_{1}+\alpha_{3}+\alpha_{4})$. Since $\alpha_{2}\in R^{x}$, $\gamma$ cannot acts as $-id$ on $R^{x}$. Hence $R^{\langle\gamma,x\rangle}\neq 0$. Suppose $|H|=2$. Since $R^{H}=\{0\}$, hence it acts as $-id$ on $R$. Then $R^{\#}$ contains roots. To conclude, there is no $p$-root pair of type $D_{4}$ if $p\geq 5$. ∎
However, we have
Lemma  5.14.
If $p=3$, $(\langle gx\rangle,D_{4})$ is a $p$-root pair.
Proof.
For $(\langle gx\rangle,D_{4})$
$R^{\#}$ $=\langle 3D_{4},2\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{1}+\alpha_{3}+\alpha_{4}\rangle$
$=\langle 3D_{4},2v_{1}-v_{2}-v_{4},v_{1}-v_{2}+2v_{3}\rangle$
hence for all elements $\lambda\in R^{\#}$, suppose that $\lambda=a(2v_{1}-v_{2}-v_{4})+b(v_{1}-v_{2}+2v_{3})+\cdots$ we have that:
[…]
If $\lambda$ is a root, then two of $\{(2a+b),(-a-b),2b,(-a)\}$ has to be 00 in $\mathbb{Z}/3\mathbb{Z}$ which is impossible. ∎
5.4. $p$-Root Pairs of Type $E$.
For lattices of type $E$, we have $\operatorname{Aut}(E_{6})=W(E_{6})\rtimes(\mathbb{Z}/2\mathbb{Z})$, $\operatorname{Aut}(E_{7})=W(E_{7})$, $\operatorname{Aut}(E_{8})=W(E_{8})$.
Lemma  5.15.
Let $(H,R)$ be a pseudo $p$-root pair, such that $|H|$ is coprime to $p$, then $S_{H}(R)$ has no roots.
Proof.
Take any $v\in S_{H}(R)$. Then $\sum\limits_{g\in H}gv=0$. Combining with $\sum\limits_{g\in H}(gv-v)\in R^{\#}$, we deduce $|H|v\in R^{\#}$. Since $\gcd(|H|,p)=1$ and $pv\in R^{\#}$, we have $v\in R^{\#}$. Therefore, $v$ is not a root. ∎
Corollary  5.16.
Suppose $(H,R)$ is a pseudo $p$-root pair, $\gcd(|H|,p)=1$ and $H\leq W(R)$. Then $(H,S_{H}(R))$ is a Leech pair.
Proof.
Since the induced action of $W(R)$ on the discriminant group $A_{R}$ is trivial, by Lemma 2.11, we know that the induced action of $H$ on the discriminant group of $S_{H}(R)$ is trivial. Then the corollary follows from Lemma 5.15. ∎
Proposition  5.17.
Let $(H,R)$ be a pseudo $p$-root pair and $R$ is of type $E_{n}$ where $n=6,7,8$, then $H\cap W(R)$ is a $p$-group.
Proof.
Suppose not, we choose a $g_{0}\in H\cap W(R)$, of order $q$ an odd prime different from $p$. We put $H_{0}=\langle g_{0}\rangle$. By Corollary 5.16, $(H_{0},S_{H_{0}}(R))$ is a Leech pair. Apparently $\operatorname{rank}(S_{H_{0}}(R))+\ell(A_{S_{H_{0}}(R)})\leq 24$, so $(H_{0},S_{H_{0}}(R))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. However, from Höhn-Mason list, this is impossible because $\operatorname{rank}(S_{H_{0}}(R))<\operatorname{rank}(R)\leq 8$. ∎
Lemma  5.18.
Suppose $(H,R)$ is a $p$-root pair with $R$ of type $E_{n}$ ($n=6,7$ or $8$), then either $p=3$ and $R=E_{6}$ or $E_{8}$, or $p=5$ and $R=E_{8}$.
Proof.
We have $\left|W(E_{8})\right|=696729600=2^{14}\cdot 3^{5}\cdot 5^{2}\cdot 7$, $\left|W(E_{7})\right|=2903040=2^{10}\cdot 3^{4}\cdot 5\cdot 7$, $\left|W(E_{6})\right|=51840=2^{7}\cdot 3^{4}\cdot 5$. Suppose $(H,R)$ is a $p$-root pair. Then for $R=E_{7},E_{8}$ only $p=3,5,7$ are possible.
For $R=E_{6}$, We first show that $H\cap W(E_{6})\neq\{id\}$. Otherwise $H\cap W(E_{6})=\{id\}$, then $|H|=2$, which implies $S_{H}(E_{6})$ is rootless by Lemma 5.15. So $R^{H}$ is nonzero, contradiction. By Proposition 5.17, $|H\cap W(E_{6})|$ is a power of $p$. Hence $p\big{|}|W(E_{6})|$, $p=3$ or $5$.
If $p=7$, then $R=E_{8}$ or $E_{7}$. The Sylow $7$-subgroup of $W(R)$ is of order $7$, hence $H\cong\mathbb{Z}/7\mathbb{Z}$. Since there are only two irreducible representation of $\mathbb{Z}/7\mathbb{Z}$ over $\mathbb{Q}$, which are of dimension $1$ and $6$, there must exist nonzero invariant vector in $R\otimes\mathbb{Q}$, contradiction.
If $p=5$, we exclude the possibilities $R=E_{6},E_{7}$ as follows. Suppose $(H,R)$ is $p$-root pair with $R=E_{6}$ or $E_{7}$. If $H\leq W(R)$, then $H$ is isomorphic to $\mathbb{Z}/5\mathbb{Z}$. Consider representations of $H$ over $\mathbb{Q}$, we know that $R^{H}$ is nonzero and hence $(H,R)$ cannot be a $p$-root pair.
If $H$ is not a subgroup of $W(R)$, then we must have $R=E_{6}$, and the order of $H$ has to be $10$. We put $H^{\prime}=H\cap W(E_{6})$, then $H^{\prime}$ is a normal subgroup of $H$ of order $5$. The representation $(E_{6})_{\mathbb{Q}}$ of $H^{\prime}$ must be a direct sum of a $2$-dimensional invariant subspace and a $4$-dimensional irreducible representation. Thus $\operatorname{rank}(E_{6}^{H^{\prime}})=2$. Next we show that $E_{6}^{H^{\prime}}$ is not rootless. Since by Sylow theory, every subgroup of $W(E_{6})$ with order $5$ is conjugate to $H^{\prime}$, we only need to find one automorphism of $E_{6}$ with order $5$ and fixing one root. Since the Dynkin graph of $A_{4}\oplus A_{1}$ is a subgraph of the Dynkin graph of $E_{6}$, there exists a primitive embedding $A_{4}\oplus A_{1}\hookrightarrow E_{6}$. By Lemma 2.12, any faithful action of $\mathbb{Z}/5\mathbb{Z}$ on $A_{4}$ extends to an action on $E_{6}$ which fixes $A_{1}$. Therefore, there exists an automorphism of $E_{6}$ with order $5$ and fixing at least one root.
We take an involution $g\in H$. Then $gH^{\prime}g^{-1}=H^{\prime}$, and $g(E_{6}^{H^{\prime}})=E_{6}^{H^{\prime}}$. Since $E_{6}^{H}=\{0\}$, the involution $g$ is a scalar by $-1$ on $E_{6}^{H^{\prime}}$. Take a root $v\in E_{6}^{H^{\prime}}$. Then $2v=-g(v)+v\in R^{\#}$, which implies $v\in R^{\#}$, contradiction. We conclude that for $5$-root pair $(H,R)$ of $E$-type, $R$ must be $E_{8}$.
If $p=3$, we show that $R\neq E_{7}$. Actually, since the exponents of $3$ in $W(E_{6})$ and $W(E_{7})$ are the same, any action of $3$-group on $E_{7}$ comes from an action on $E_{6}\subset E_{7}$. This implies the existence of a fixed vector in $E_{7}$, which is a contradiction. ∎
We can conclude that:
Proposition  5.19.
Let $(H,R)$ be a $p$-root pair. If $p\geq 7$, then $R$ has no component of type $E$.
We now analyze $p=5$ and $R=E_{8}$. Since $\left|W(E_{8})\right|=696729600=2^{14}\cdot 3^{5}\cdot 5^{2}\cdot 7$, the Sylow-$5$ subgroup of $W(E_{8})$ is of order $25$. The following description of $E_{8}$ lattice is referred to Kac [Kac90, §6.7]:
  1. (1)
    $E_{8}=\{\sum k_{i}v_{i}\in\mathbb{R}^{8}|k_{i}\in\mathbb{Z}\;\text{or all}\;k_{i}\in 1/2\mathbb{Z},\text{and}\;\sum k_{i}\in 2\mathbb{Z}\}$,
  2. (2)
    roots:{±$\{\pm v_{i}\pm v_{j},1\leq i<j\leq 8\}\cup\{\frac{1}{2}(\pm v_{1}\pm v_{2}\pm\cdots\pm v_{8})|\text{even number of minuses}\}$,
  3. (3)
    positive simple roots: $\{\alpha_{i}=v_{i+1}-v_{i+2}|i=1,\cdots,6\}$ and $\alpha_{7}=\frac{1}{2}(v_{1}-v_{2}-\cdots-v_{7}+v_{8})$ and $\alpha_{8}=v_{7}+v_{8}$,
  4. (4)
    the highest root is $\theta=v_{1}+v_{2}$.
The roots $\alpha_{i}$ ($1\leq i\leq 8$) and $-\theta$ form the affine Dynkin graph:
[…]
By deleting the node corresponding to $\alpha_{4}$, we obtain a subgraph of $\widetilde{E}_{8}$ which is the Dynkin graph of $A_{4}\oplus A_{4}$. So we have an embedding of $A_{4}\oplus A_{4}$ into $E_{8}$. Explictly, $\{-\theta,\alpha_{1},\alpha_{2},\alpha_{3}\}$ and $\{\alpha_{8},\alpha_{5},\alpha_{6},\alpha_{7}\}$ are naturally bases for the two pieces of $A_{4}$ respectively.
For any root $\alpha$, we simply put $s_{\alpha}$ as the reflection defined by $\alpha$, i.e., $s_{\alpha}(x)=x-\langle x,\alpha\rangle\alpha$.
Lemma  5.20.
There is a faithful action of $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}\subset W(A_{4})\times W(A_{4})$ on $E_{8}$.
Proof.
We denote the natural generators of $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}\subset W(A_{4})\times W(A_{4})=S_{5}\times S_{5}$, by $a$ and $b$. We want to extend the action of $a,b$ on the lattice $A_{4}$ to an action on $E_{8}$.
We define $a(-\theta)=\alpha_{1},a(\alpha_{1})=\alpha_{2},a(\alpha_{2})=\alpha_{3},a(\alpha_{3})=\theta-\alpha_{1}-\alpha_{2}-\alpha_{3}$, and $a(\alpha_{i})=\alpha_{i}$ for $i=5,6,7,8$. Since $\theta=2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+5\alpha_{4}+6\alpha_{5}+4\alpha_{6}+2\alpha_{7}+3\alpha_{8}$ and $a(-\theta)=\alpha_{1}$, thus $a(\alpha_{4})=-\theta+\sum\limits_{i=1}^{4}\alpha_{i}$. We can verify that $a=s_{-\theta}s_{\alpha_{1}}s_{\alpha_{2}}s_{\alpha_{3}}$, hence it defines an element of $W(E_{8})$, which we still denote by $a$.
We define $b(\alpha_{i})=\alpha,i=1,2,3,b(\theta)=\theta,b(\alpha_{8})=\alpha_{5},b(\alpha_{5})=\alpha_{6},b(\alpha_{6})=\alpha_{7},b(\alpha_{7})=-\sum\limits_{i=5}^{8}\alpha_{i}$, thus $b(\alpha_{4})=\alpha_{4}+\alpha_{5}+\alpha_{8}$. We can verify that $b=s_{\alpha_{8}}s_{\alpha_{5}}s_{\alpha_{6}}s_{\alpha_{7}}$, hence it defines an element of $W(E_{8})$, which we still denote by $b$.
We can verify that $ab=ba$. For example $ab(\alpha_{4})=ba(\alpha_{4})=-\theta+\sum\limits_{i=1}^{5}\alpha_{i}+\alpha_{8}$. In particular, $\langle a,b\rangle$ is a Sylow-$5$ subgroup of $W(E_{8})$. ∎
Proposition  5.21.
If $p=5$, $(H,E_{8})$ is a $p$-root pair, then $H\cong\mathbb{Z}/5\mathbb{Z}$. Moreover, there exists a $p$-root pair $(H,E_{8})$ with $H\cong\mathbb{Z}/5\mathbb{Z}$.
Proof.
First notice that any subgroup of $W(E_{8})$ of order $25$ is a Sylow subgroup which is conjugate to the action we construct above. Since $a(\alpha_{4})-\alpha_{4}=-\theta+\alpha_{1}+\alpha_{2}+\alpha_{3}=-v_{1}-v_{5}$ is a root, $H$ cannot be a $5$-Sylow subgroup of $W(E_{8})$. The last statement follows from [GL11, Remark 3.3]. ∎
5.5. On Reducible $p$-Root Pairs
Suppose $R=\oplus R_{i}$, then we define $H_{i}$ to be the image of $\{g\in H\big{|}g(R_{i})=R_{i}\}$ in $\operatorname{Aut}(R_{i})$, which is called the stabilizer of $R_{i}$.
Lemma  5.22.
If $(H,R)$ is a pseudo $p$-root pair, then each $(H_{i},R_{i})$ is also a pseudo $p$-root pair. If $(H,R)$ is a $p$-root pair, then each $(H_{i},R_{i})$ is also a $p$-root pair.
Proof.
The first statement is clear. Suppose $(H,R)$ is a $p$-root pair, to show that $(H_{i},R_{i})$ is a $p$-root pair, we only need to show $R_{i}^{H_{i}}=\{0\}$. Otherwise there is a nonzero vector $v\in R_{i}^{H_{i}}$. The factor of $\sum\limits_{g\in H}gv$ in $R_{i}$ equals to |$|\{g\in H\big{|}g(R_{i})=R_{i}\}|v$, hence nonzero. Meanwhile $\sum\limits_{g\in H}gv\in R^{H}=\{0\}$, a contradiction. ∎
Lemma  5.23.
If $(H,R)$ is a pseudo $p$-root pair, then $H\cap W(R)$ is a $p$-group.
Proof.
Let $H_{i}$ be as before. Then we have a natural map:
$H\cap W(R)\rightarrow\prod H_{i}\cap W(R_{i})$
since $H\cap W(R)$ stabilizes each irreducible component of $R$. It is also injective because if an element stabilizes and acts trivially on each irreducible component then it acts trivially on $R$. Then the lemma follows from Lemma 5.2, Proposition 5.9 and Proposition 5.17. ∎
6. Application of $p$-Root Pairs
Let $G$ be a finite group acting faithfully and symplectically on a supersingular K3 surface $X$ with Artin invariant $\sigma\geq 2$. Recall that the action of $G$ on $N_{p,\sigma}$ can be extended to $N_{p,1}$ since $G$ acts trivially on $A_{N_{p,\sigma}}$. This gives rise to $(G_{R},R)$ where $R\subset S_{G}(N_{p,1})$ is the root lattice. Moreover we have the following exact sequence:
$1\rightarrow G_{L}\rightarrow G\rightarrow G_{R}\rightarrow 1$
And by Lemma 3.16, $(G_{R},R)$ is a $p$-root pair. In this section, we apply our results on $p$-root pairs to show that when $p\geq 13$, we always have $(p,|G|)=1$, which recovers Dolgachev–Keum [DK09, Theorem 2.1] for supersingular K3. And then we calculate the upper bounds of the $p$-part of $G$ for $p=3,5,7,11$.
6.1. New Proof for Dolgachev–Keum’s Result
We start by considering a $p$-root pair $(H,R)$ with $R=A_{p-1}$.
Definition  6.1.
For $R=A_{p-1}$ with chosen simple roots $\alpha_{i}=v_{i}-v_{i+1}$, we define $T(R)=\{\sum\limits_{i=1}^{p-1}k_{i}\alpha_{i}\big{|}k_{i}\in\mathbb{Z},\sum\limits_{i=1}^{p-1}k_{i}\equiv 0\;(\mathrm{mod}\;p)\}$.
Lemma  6.2.
We have
$T(A_{p-1})\subset A_{p-1}\subset A_{p-1}^{\vee}\subset T(A_{p-1})^{\vee}$
with each adjacent inclusion of index $p$. And
  1. (1)
[…]
  2. (2)
    $q_{T(A_{p-1})}=p^{+3}$.
Proof.
Take $\omega_{i}\in A_{p-1}^{\vee}$ with $\omega_{i}(\alpha_{j})=\delta_{ij}$, $11\leq i,j\leq p-$ 1. We can also regard $\omega_{i}$ as elements in $(A_{p-1})_{\mathbb{Q}}$. Then $A_{p-1}^{\vee}=\mathrm{Span}_{\mathbb{Z}}\{\omega_{1},\cdots,\omega_{p-1}\}$. We have
$\omega_{1}=v_{1}-{1\over p}(v_{1}+\cdots+v_{p})={1\over p}(\alpha_{p-1}+2\alpha_{p-2}+\cdots+(p-1)\alpha_{1}).$
Then the discriminant group of $A_{p-1}$ is a group of order $p$ with a generator $[\omega_{1}]$. The value of the finite quadratic form $q_{A_{p-1}}$ at $[\omega_{1}]$ is $1-{1\over p}\in\mathbb{Q}/2\mathbb{Z}$. This shows (1).
For (2) it suffices to show $pT(A_{p-1})^{\vee}\subset T(A_{p-1})$. We have
$T(A_{p-1})^{\vee}=\mathrm{Span}_{\mathbb{Z}}(\omega_{1},\cdots,\omega_{p-1},{1\over p}\sum\limits_{i=1}^{p-1}\omega_{i}).$
Since $p\omega_{1}=-(\alpha_{p-1}+2\alpha_{p-2}+\cdots+(p-1)\alpha_{1})\in T(A_{p-1})$, we have $pA_{p-1}^{\vee}\subset T(A_{p-1})$. We can directly check that
$\sum\limits_{i=1}^{p-1}\omega_{i}=\sum\limits_{i=1}^{p-1}i(p-i)\alpha_{i}\in T(A_{p-1}).$
The Lemma then follows. ∎
Lemma  6.3.
Suppose $(H,R)$ is a $p$-root pair with $R=A_{p-1}$. Then $H\cap W(R)$ has order $p$. Take a generator $g\in H\cap W(R)$, then it is a cycle in $S_{p}$ of length $p$. Suppose simple roots $\alpha_{i}=v_{i}-v_{i+1}$ satisfy $g(v_{i})=v_{i+1}$ for $11\leq i\leq p-$ 1 and $g(v_{p})=v_{1}$. Then $R^{\#}=T(A_{p-1})$. And the induced action of $H$ on the discriminant group of $T(A_{p-1})$ is nontrivial.
Proof.
The sublattice $T(A_{p-1})$ is generated by $\alpha_{1}-\alpha_{2},\cdots,\alpha_{p-2}-\alpha_{p-1},pA_{p-1}$. Since
$g(\alpha_{i})-\alpha_{i}=\alpha_{i+1}-\alpha_{i},1\leq i\leq p-2,g(\alpha_{p-1})-\alpha_{p-1}=-\sum_{j=1}^{p-1}\alpha_{j}-\alpha_{p-1}$
we have $T_{A_{p-1}}\subset R^{\#}$. Because $[A_{p-1}:T(A_{p-1})]=p$ and $R^{\#}$ is strictly contained in $R=A_{p-1}$, we must have $T_{A_{p-1}}=R^{\#}$.
We have ${1\over p}\sum\limits_{i=1}^{p-1}\omega_{i}\in(T(A_{p-1}))^{\vee}$. We can check that:
(3) $g({1\over p}\sum\limits_{i=1}^{p-1}\omega_{i})-{1\over p}\sum\limits_{i=1}^{p-1}\omega_{i}=-\omega_{1}\notin A_{p-1}.$
Hence $H$ acts nontrivially on the discriminant group of $T(A_{p-1})$. ∎
Suppose $(H,R)$ is a $p$-root pair with $R=A_{p-1}^{k}$, where $k$ is a positive integer. We write $R=R_{1}\oplus\cdots\oplus R_{k}$ with each $R_{i}$ isomorphic to $A_{p-1}$. Then $(H_{i},R_{i})$ is a $p$-root pair by Lemma 5.22. By Lemma 6.3, $H_{i}\cap W(R_{i})$ contains a generator of order $p$, which gives rise to a sublattice $T(R_{i})\subset R_{i}$. Then we have the following lemma:
Lemma  6.4.
Under the above setup, we have $\oplus_{i=1}^{k}T(R_{i})\subset R^{\#}$, and for any even lattice $M$ with $\oplus_{i=1}^{k}T(R_{i})\subset M\subsetneq R$ which is preserved by the action of $H$, the induced action of $H$ on $A_{M}$ is nontrivial.
Proof.
By Lemma 6.3, we have $T(R_{i})=R_{i}^{\#}$ with respect to the $p$-root pair $(H_{i},R_{i})$. Therefore, $R^{\#}$ contains $\oplus_{i=1}^{k}T(R_{i})$ as a sublattice. Now suppose $M$ to be a lattice with $\oplus_{i=1}^{k}T(R_{i})\subset M\subsetneq R=\oplus_{i=1}^{k}R_{i}$ and is preserved by the action of $H$. We have
$\oplus_{i=1}^{k}T(R_{i})\subset M\subsetneq\oplus_{i=1}^{k}R_{i}\subsetneq\oplus_{i=1}^{k}R_{i}^{\vee}\subsetneq M^{\vee}\subset\oplus_{i=1}^{k}T(R_{i})^{\vee}.$
We can take an element $x=(x_{1},\cdots,x_{k})\in M^{\vee}$, such that $x_{1}\notin R_{1}^{\vee}$. Take $h$ a generator of $H_{1}\cap W(R_{1})\leq H$. From Equation (3), we know $h(x_{1})-x_{1}\notin R_{1}$. Therefore, $h(x)-x\notin R$. This implies that the action of $h$ on $A_{M}$ is nontrivial. ∎
The following theorem is obtained by Dolgachev–Keum [DK09, Theorem 2.1] through a geometric argument using fixed loci and fibrations. For supersingular cases, we provide a lattice theoretic proof.
Theorem  6.5 (Dolgachev–Keum).
Let $X$ be a supersingular K3 surface. Then if $p\geq 13$, for any finite group $G$ acting symplectically and faithfully on $X$, we have $(p,|G|)=1$.
Proof.
When $\sigma(X)=1$, this follows from the classfication result in [OS24]. We now deal with the cases that $\sigma(X)\geq 2$. We argue by contradiction.
Consider the sequence (2). By Proposition 3.15, we have shown that $(G_{L},S_{G_{L}}(N_{p,1}))$ is a Höhn–Mason pair, then by the Höhn–Mason list, $G_{L}$ is tame. Hence $G\neq G_{L}$, and $R$ is nontrivial.
Without loss of generality, we assume that $G\cong\mathbb{Z}/p\mathbb{Z}$. By Proposition 5.11, there is no component of type D. By Lemma 5.19, there is no component of type E. Then all the possible $R$ is of type $A$. Again by Lemma 5.22 and Proposition 5.5, $R$ has to be irreducible. Hence $R=A_{p-1}$.
Suppose $\operatorname{rank}(S_{G}(N_{p,\sigma}))=\operatorname{rank}(A_{p-1})$. Since $A_{A_{p-1}}\cong\mathbb{Z}/p\mathbb{Z}$ and $S_{G}(N_{p,\sigma})$ is rootless, then $S_{G}(N_{p,\sigma})=T(A_{p-1})$. By Lemma 6.4, the action of $G_{R}$ on $A_{S_{G}(N_{p,\sigma})}$ is nontrivial, contradiction. Therefore, $\operatorname{rank}(S_{G}(N_{p,\sigma}))>\operatorname{rank}(A_{p-1})$. The action of $g$ on $S_{G}(N_{p,\sigma})$ has no fixed vector, which implies $(p-1)\big{|}\operatorname{rank}(S_{G}(N_{p,\sigma}))$. Then $\operatorname{rank}(S_{G}(N_{p,\sigma}))\geq 2(p-1)\geq 24$, contradiction. ∎
6.2. Upper Bounds of Wild Degrees
We denote the exponent of $p$ in $|G|$ by $\nu_{p}(G)$, and call it the wild degree of $G$. In this subsection we aim to give an upper bound of the wild degree for a finite group $G$ of symplectic automophisms of a supersingular K3 surface in characteristic $p=11,7,5,3$. We start from the following lemma which specializes results on $p$-root pairs to $R\subset S_{G}(N_{p,1})$.
Lemma  6.6.
  1. (1)
    For $p=11$, the only possible irreducible component of $R$ is $A_{10}$.
  2. (2)
    For $p=7$, the only possible irreducible component of $R$ is $A_{6}$.
  3. (3)
    For $p=5$, all possible irreducible components of $R$ are: $A_{4},E_{8}$.
  4. (4)
    For $p=3$, all possible irreducible components of $R$ are: $A_{2},A_{8},D_{4},E_{6},E_{8}$.
Proof.
For $p=11,7$, by Proposition 5.11, 5.13 and 5.19, all components of $R$ are of type A. Since $\operatorname{rank}R\leq 22$, by Lemma 5.22 and Proposition 5.5, each component of $R$ is isomorphic to $A_{p-1}$.
For $p=5$, by Proposition 5.11, 5.13, there is no irreducible component of type $D$. Then the conclusion follows from Lemma 5.22, Proposition 5.5, 5.21.
For $p=3$, by Proposition 5.11, Lemma 5.14, $D_{4}$ is the only possibility of type D. The remaining conclusion follows from Lemma 5.18, 5.22 and Proposition 5.5. ∎
Proposition  6.7.
For $p=11$, we have $\nu_{p}(G)\leq 1$.
Proof.
By Lemma 6.6, $R\cong A_{10}$ or $R\cong A_{10}\oplus A_{10}$. By Lemma 5.2, $\nu_{p}(G_{R})\leq 2$.
Suppose $\nu_{p}(G_{R})=2$, then $R=R_{1}\oplus R_{2}$ with $R_{1}\cong R_{2}\cong A_{10}$ and $G_{R}\cap W(R)=(\mathbb{Z}/11\mathbb{Z})^{2}$. Take generators $g_{1},g_{2}$ for $G_{R}\cap W(R)$, such that $g_{1}$ acts trivially on $R_{2}$, and $g_{2}$ acts trivially on $R_{1}$. Let $\alpha_{i}=v_{i}-v_{i+1}$ ($1\leq i\leq 10$) be simple roots for $R_{1}$, such that $g_{1}(v_{i})=v_{i+1}$ for $1\leq i\leq 10$ and $g_{1}(v_{11})=v_{1}$. Then we can define $T(R_{1})$. Similarly we can define $T(R_{2})$. Then $R^{\#}=T(R_{1})\oplus T(R_{2})$.
The action of $g$ on $S_{G}(N_{p,\sigma})$ has no fixed vector, which implies $(p-1)\big{|}\operatorname{rank}(S_{G}(N_{p,\sigma})$. We must have $\operatorname{rank}(S_{G}(N_{p,1}))=20$. Since $q_{R}=11^{+2}$, $R$ cannot have nontrivial saturation. Thus $S_{G}(N_{p,1})=R$. Hence we have:
$R^{\#}\subset S_{G}(N_{p,\sigma})\subsetneq R.$
However by Lemma 6.4, the action of $G_{R}$ on $A_{S_{G}(N_{p,\sigma})}$ is nontrivial, contradiction. Therefore, $\nu_{p}(G_{R})\leq 1$.
By Proposition 2.9, the Leech pair $(G_{L},S_{G_{L}}(N_{p,1}))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Since $G_{L}$ acts trivially on $R$, by the rank $\geq 10$ cases of Höhn-Mason [HM16, Table 1], we know that $|G_{L}|$ is coprime to $p$. Hence we can conclude that $\nu_{p}(G)\leq 1$. ∎
Proposition  6.8.
We have the following upper bounds of the wild degree of $G$:
  1. (1)
    For $p=7$, $\nu_{p}(G)\leq 3$.
  2. (2)
    For $p=5$, $\nu_{p}(G)\leq 6$.
  3. (3)
    For $p=3$, $\nu_{p}(G)\leq 14$.
Proof.
Suppose $p=7$. By Lemma 6.6,
$R\cong A_{6},\;\text{or}\;R\cong A_{6}\oplus A_{6},\;\text{or}\;R\cong A_{6}\oplus A_{6}\oplus A_{6}$
If $R\cong A_{6}$, then $\nu_{p}(G_{R})\leq 1$. By Proposition 3.15, the Leech pair $(G_{L},S_{G_{L}}(N_{p,1}))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Since $G_{L}$ acts trivially on $R$, by the rank $\leq 18$ cases of Table LABEL:table:_sublist_of_HM, we know that the $\nu_{p}(G_{L})\leq 1$ (the only case with $\nu_{p}(G_{L})=1$ is case 52). Hence we can conclude that $\nu_{p}(G)\leq 2$ in this case.
If $R\cong A_{6}\oplus A_{6}$, then $\nu_{p}(G_{R})\leq 2$. By Proposition 3.15, the Leech pair $(G_{L},S_{G_{L}}(N_{p,1}))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. Since $G_{L}$ acts trivially on $R$, by the rank $\leq 12$ cases of Table LABEL:table:_sublist_of_HM, we know that $|G_{L}|$ is coprime to $p$. Hence we can conclude that $\nu_{p}(G)\leq 2$ in this case.
If $R\cong A_{6}\oplus A_{6}\oplus A_{6}$, $\nu_{p}(G_{R})\leq 3$. Again, $G_{L}$ acts trivially on $R$, by the rank $\leq 6$ cases of Table LABEL:table:_sublist_of_HM, we know that $|G_{L}|$ is coprime to $p$. Hence $\nu_{p}(G)\leq 3$ in this case.
Suppose $p=5$. By Lemma 6.6, there are two possibilities, components of $R$ can be of type $A_{4}$ or $E_{8}$. By Proposition 5.21, each $E_{8}$ component contributes at most one copy of $\mathbb{Z}/5\mathbb{Z}$. By Proposition 2.9, the Leech pair $(G_{L},S_{G_{L}}(N_{p,1}))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$. By the $\operatorname{rank}\leq 16$ cases in Table LABEL:table:_sublist_of_HM, $|G_{L}|$ is coprime to $p$. Hence we can conclude that if there exist $E_{8}$ components, $\nu_{p}(G)\leq 4$ and the equality holds only if $R=E_{8}\oplus A_{4}\oplus A_{4}\oplus A_{4}$.
Now we assume that there is no $E_{8}$ component. Then as in $p=7$, we can analyze $\nu_{p}(G_{R}),\mu_{p}(G_{L})$ by the number of $A_{4}$ components in $R$ and Table LABEL:table:_sublist_of_HM. To reach the maximal $\nu_{p}$, $R\cong A_{4}\oplus A_{4}\oplus A_{4}\oplus A_{4}\oplus A_{4}$. Considering the possible extra $\mathbb{Z}/5\mathbb{Z}$ given by the permutation, we have $\nu_{p}(G)\leq 6$.
Finally suppose $p=3$. By Lemma 6.6, all possible irreducible components of $R$ are $A_{2},A_{8},D_{4},E_{6},E_{8}$. We also have
  1. (1)
    If $(G,A_{2})$ is a $p$-root pair, $\nu_{p}(G)\leq 1$.
  2. (2)
    If $(G,A_{8})$ is a $p$-root pair, $\nu_{p}(G)\leq 4$.
  3. (3)
    If $(G,D_{4})$ is a $p$-root pair, $\nu_{p}(G)\leq 1$ (see Lemma 5.14).
  4. (4)
    If $(G,E_{6})$ is a $p$-root pair, $\nu_{p}(G)\leq 4$.
  5. (5)
    If $(G,E_{8})$ is a $p$-root pair, $\nu_{p}(G)\leq 5$.
Compare $(G,A_{2}\oplus A_{2}\oplus A_{2}\oplus A_{2})$ with $(G,A_{8})$, there can be an extra $S_{4}$ acting on $A_{2}\oplus A_{2}\oplus A_{2}\oplus A_{2}$. Hence to reach the maximal possible wild degree, we only need to consider $A_{2},E_{6},E_{8}$ components.
  1. (1)
    If $\operatorname{rank}R=2$, $R=A_{2}$, then $\nu_{p}(G_{R})\leq 1$. By rank $\leq 22$ cases in Table LABEL:table:_sublist_of_HM, $\nu_{p}(G_{L})\leq 6$ (see case 101). Hence $\nu_{p}(G)\leq 7$.
  2. (2)
    If $\operatorname{rank}R=4$, we can choose $R=A_{2}\oplus A_{2}$, then $\nu_{p}(G_{R})\leq 2$. By rank $\leq 20$ cases in Table LABEL:table:_sublist_of_HM, $\nu_{p}(G_{L})\leq 6$ (see case 101). Hence $\nu_{p}(G)\leq 8$.
  3. (3)
    If $\operatorname{rank}R=6$, we have $\nu_{p}(G_{R})\leq 4$. By rank $\leq 18$ cases in Table LABEL:table:_sublist_of_HM, $\nu_{p}(G_{L})\leq 5$ (see case 35). Hence $\nu_{p}(G)\leq 9$.
  4. (4)
    If $\operatorname{rank}R=8$, we have $\nu_{p}(G_{R})\leq 5$. By rank $\leq 16$ cases in Table LABEL:table:_sublist_of_HM, $\nu_{p}(G_{L})\leq 1$. Hence $\nu_{p}(G)\leq 6$.
  5. (5)
    If $\operatorname{rank}R=10$, we have $\nu_{p}(G_{R})\leq 6$ ($R=A_{2}^{\oplus 5}$). By rank $\leq 14$ cases in Table LABEL:table:_sublist_of_HM, $\nu_{p}(G_{L})\leq 1$. Hence $\nu_{p}(G)\leq 7$.
  6. (6)
    If $\operatorname{rank}R=12$, we have $\nu_{p}(G_{R})\leq 8$. By rank $\leq 12$ cases in Table LABEL:table:_sublist_of_HM, $\nu_{p}(G_{L})\leq 1$. Hence $\nu_{p}(G)\leq 9$.
  7. (7)
    If $\operatorname{rank}R\geq 14$, $\nu_{p}(G_{R})\leq 14$, the equality holds when $R=A_{2}^{\oplus 10}$. By rank $\leq 10$ cases in Table LABEL:table:_sublist_of_HM, $\nu_{p}(G_{L})=0$. Hence $\nu_{p}(G)\leq 14$.
To conclude, we have $\nu_{p}(G)\leq 14$. ∎
Appendix A Höhn–Mason Pairs Admitting Primitive Embedding into $N_{p,1}$
For a Leech pair $(G,S)$ from Höhn–Mason list, we need to justify for which odd prime $p$ is $S$ a primitive sublattice of $N_{p,1}$.
[…]
Table 1. Höhn–Mason Pairs Admitting Primitive Embedding into $N_{p,1}$
No. rk(S) Group Order $q_{S}$ p
$1$ 00 $1$ $1$ $1$ any
$2$ $8$ $2$ $2$ $2_{\operatorname{II}}^{+8}$ any
$3$ $12$ $2^{2}$ $4$ $2_{\operatorname{II}}^{-6}4_{\operatorname{II}}^{-2}$ any
$4$ $12$ $3$ $3$ $3^{+6}$ any
$6$ $14$ $2^{3}$ $8$ $2_{\operatorname{II}}^{+6}4_{2}^{+2}$ any
$7$ $14$ $S_{3}$ $6$ $2_{\operatorname{II}}^{-2}3^{+5}$ any
$9$ $14$ $4$ $4$ $2_{2}^{+2}4_{\operatorname{II}}^{+4}$ any
$10$ $15$ $2^{4}$ $16$ $2_{\operatorname{II}}^{+6}8_{1}^{+1}$ any
$13$ $15$ $D_{8}$ $8$ $4_{1}^{+5}$ any
$15$ $16$ $A_{3,3}$ $18$ $3^{+4}9^{-1}$ any
$17$ $16$ $\Gamma_{2}a_{1}=2\times D_{8}$ $16$ $2_{\operatorname{II}}^{+2}4_{0}^{+4}$ any
$18$ $16$ $D_{12}$ $12$ $2_{\operatorname{II}}^{+4}3^{+4}$ any
$19$ $16$ $A_{4}$ $12$ $2_{\operatorname{II}}^{-2}4_{\operatorname{II}}^{-2}3^{+2}$ any
$20$ $16$ $D_{10}$ $10$ $5^{+4}$ any
$24$ $17$ $32^{4}:$ 3 $48$ $2_{\operatorname{II}}^{-4}8_{1}^{+1}3^{-1}$ any
[…]
$28$ $17$ $\Gamma_{5}a_{1}=Q_{8}*Q_{8}$ $32$ $4_{7}^{+5}$ any
$29$ $17$ $S_{4}$ $24$ $4_{3}^{+3}3^{+2}$ any
$31$ $17$ $Q_{8}$ $8$ $2_{7}^{-3}8_{\operatorname{II}}^{-2}$ any
$35$ $18$ $23^{1+4}:$ 2 $486$ $3^{+5}$ $3$
$37$ $18$ $4^{2}.A_{4}$ $192$ $2_{\operatorname{II}}^{-2}8_{6}^{-2}$ any
$38$ $18$ $2^{4}.D_{6}$ $96$ $2_{\operatorname{II}}^{-2}4_{7}^{+1}8_{1}^{+1}3^{-1}$ any
$39$ $18$ $A_{4,3}$ $72$ $4_{\operatorname{II}}^{-2}3^{-3}$ any
$43$ $18$ $\Gamma_{25}a_{1}$ $64$ $4_{5}^{+3}8_{1}^{+1}$ any
$44$ $18$ $A_{5}$ $60$ $2_{\operatorname{II}}^{-2}3^{+1}5^{-2}$ any
$45$ $18$ $2\times S_{4}$ $48$ $2_{\operatorname{II}}^{+2}4_{2}^{+2}3^{+2}$ any
$46$ $18$ $43^{2}.$4 $36$ $2_{6}^{-2}3^{+2}9^{-1}$ any
$47$ $18$ $S_{3,3}$ $36$ $2_{\operatorname{II}}^{-2}3^{+3}9^{-1}$ any
$52$ $18$ $F_{21}$ $21$ $7^{+3}$ any
$53$ $18$ $\mathrm{Hol}_{4}$ $20$ $2_{6}^{-2}5^{+3}$ any
$55$ $18$ $\Gamma_{3}a_{2}=\mathrm{SD}_{16}$ $16$ $2_{7}^{+1}4_{7}^{+1}8_{\operatorname{II}}^{+2}$ any
$68$ $19$ $2.23^{1+4}:$ 2.2 $972$ $2_{1}^{+1}3^{-4}$ 3
$69$ $19$ $M_{20}$ $960$ $2_{\operatorname{II}}^{-2}8_{1}^{+1}5^{-1}$ any
$70$ $19$ $4^{2}.S_{4}$ $384$ $4_{7}^{+1}8_{6}^{+2}$ any
$72$ $19$ $A_{6}$ $360$ $4_{5}^{-1}3^{+2}5^{+1}$ any
$73$ $19$ $A_{4,4}$ $288$ $2_{\operatorname{II}}^{+2}8_{1}^{+1}3^{+2}$ any
$74$ $19$ $2^{4}:D_{12}$ $192$ $4_{2}^{-2}8_{1}^{+1}3^{-1}$ any
$76$ $19$ $(Q_{8}*Q_{8}):S_{3}$ $192$ $4_{7}^{-3}3^{+1}$ any
$77$ $19$ $L_{2}(7)$ $168$ $4_{1}^{+1}7^{+2}$ any
$82$ $19$ $S_{5}$ $120$ $4_{3}^{-1}3^{+1}5^{-2}$ any
$84$ $19$ $M_{9}$ $72$ $2_{7}^{-3}3^{-1}9^{-1}$ any
$85$ $19$ $3^{2}.D_{8}$ $72$ $4_{1}^{+1}3^{+2}9^{-1}$ any
$87$ $19$ $T_{48}$ $48$ $2_{7}^{+1}8_{\operatorname{II}}^{-2}3^{-1}$ any
$101$ $20$ $3^{4}:A_{6}$ $29160$ $3^{+2}9^{+1}$ $3$
$102$ $20$ $L_{3}(4)=M_{21}$ $20160$ $2_{\operatorname{II}}^{-2}3^{-1}7^{-1}$ $3,7$ or $({21\over p})=-1$
$106$ $20$ $2^{4}:A_{6}$ $5760$ $4_{5}^{-1}8_{1}^{+1}3^{+1}$ $3$ or $({6\over p})=-1$
$108$ $20$ $A_{7}$ $2520$ $3^{+1}5^{+1}7^{+1}$ $3,5,7$ or $({105\over p})=-1$
$109$ $20$ $3^{1+4}:2.2^{2}$ $1944$ $2_{2}^{+2}3^{+3}$ $3$
$110$ $20$ $22^{4}:S_{5}=M_{20}:$ 2 $1920$ $4_{3}^{-1}8_{1}^{+1}5^{-1}$ $5$ or $({10\over p})=-1$
$111$ $20$ $2^{3}:L_{2}(7)$ $1344$ $4_{2}^{+2}7^{+1}$ $7$ or $({7\over p})=-1$
$112$ $20$ $Q(3^{2}:2)=2^{2}.A_{4,4}$ $1152$ $8_{6}^{-2}3^{-1}$ $3$ or $({3\over p})=-1$
$118$ $20$ $S_{6}$ $720$ $2_{\operatorname{II}}^{-2}3^{+2}5^{+1}$ $3,5$ or $({5\over p})=-1$
$119$ $20$ $M_{10}$ $720$ $2_{3}^{-1}4_{7}^{+1}3^{-1}5^{+1}$ $3,5$ or $({30\over p})=-1$
$120$ $20$ $L_{2}(11)$ $660$ $11^{+2}$ $11$
$121$ $20$ $2^{4}:(S_{3}\times S_{3})$ $576$ $4_{7}^{+1}8_{1}^{+1}3^{+2}$ $({2\over p})=-1$
$122$ $20$ $45^{1+2}:$ 4 $500$ $5^{+3}$ 5
$128$ $20$ $(3\times A_{5}):2$ $360$ $3^{-2}5^{-2}$ $5$
$129$ $20$ $2\times L_{2}(7)$ $336$ $2_{\operatorname{II}}^{+2}7^{+2}$ $7$
$134$ $20$ $3^{2}:\mathrm{SD}_{16}$ $144$ $2_{1}^{+1}4_{1}^{+1}3^{-1}9^{-1}$ $({6\over p})=-1$
$163$ $21$ $U_{4}(3)$ $3265920$ $4_{7}^{+1}3^{+2}$ $3$
$165$ $21$ $M_{22}$ $443520$ $4_{5}^{-1}11^{+1}$ $11$
$167$ $21$ $U_{3}(5)$ $126000$ $2_{7}^{+1}5^{-2}$ $5$
[…]
$172$ $21$ $2^{4}.A_{7}=2^{4}:A_{7}$ $40320$ $8_{1}^{+1}7^{+1}$ $7$
$175$ $21$ $A_{8}$ $20160$ $4_{1}^{+1}3^{+1}5^{+1}$ $5$
$182$ $21$ $M_{11}$ $7920$ $2_{7}^{+1}3^{-1}11^{-1}$ $11$
$183$ $21$ $[2^{4}3].S_{5}=2^{4}:(3\times S_{5})$ $5760$ $8_{1}^{+1}3^{-1}5^{-1}$ $5$
Remark  A.1.
There are some typos about $2$-parts in discriminants in Höhn–Mason. We corrects the typos we found by comparing [HM16] to [Has12]. For case 31, $2_{3}^{+3}8_{\operatorname{II}}^{-2}$ is corrected to $2_{7}^{-3}8_{\operatorname{II}}^{-2}$. For case 46, $2_{2}^{+2}3^{+2}9^{-1}$ is corrected to $2_{6}^{-2}3^{+2}9^{-1}$. For case 53, $2_{2}^{+2}5^{+3}$ is corrected to $2_{6}^{-2}5^{+3}$. For case 55, $2_{5}^{+1}4_{1}^{+1}8_{\operatorname{II}}^{+2}$ is corrected to $2_{7}^{+1}4_{7}^{+1}8_{\operatorname{II}}^{+2}$. For case 70, $4_{3}^{+1}8_{2}^{+2}$ is corrected to $4_{7}^{+1}8_{6}^{+2}$. For case 84, $2_{2}^{+3}3^{-1}9^{-1}$ is corrected to $2_{7}^{-3}3^{-1}9^{-1}$. For case 119, $2_{5}^{+1}4_{1}^{+1}3^{-1}5^{+1}$ is corrected to $2_{3}^{-1}4_{7}^{+1}3^{-1}5^{+1}$.
Combining with Kondō [Kon06], in the case $(G,q_{S})\cong([2^{4}3].S_{5},8_{1}^{+1}3^{-1}5^{-1})$, we actually have $G\cong 2^{4}:(3\times S_{5})$.

### v3

@@
 Abstract.
-In 2009, Dolgachev–Keum show that finite groups of tame symplectic automorphisms of K3 surfaces in positive characteristics are subgroups of the Mathieu group of degree $23$ under mild conditions.
+In 2009, Dolgachev–Keum showed that finite groups of tame symplectic automorphisms of K3 surfaces in positive characteristics are subgroups of the Mathieu group of degree $23$.
 In this paper, we utilize lattice-theoretic methods to investigate symplectic actions of finite groups $G$ on K3 surfaces in odd characteristics.
-In the tame cases (i.e., the order of $G$ is coprime with $p$) and all the superspecial cases, one can associate with the action a Leech pair which can be detected via Höhn–Mason’s list.
-Notice that the superspecial cases have been resolved by Ohashi–Schütt.
-If the K3 surface is supersingular with Artin invariant at least two, we develop a new machinery called $p$-root pairs to detect possible symplectic finite group actions (without the assumption of tameness).
+For supersingular K3 surfaces with Artin invariants at least two, we develop a new machinery called $p$-root pairs to constrain possible symplectic finite group actions (without the assumption of tameness).
 The concept of $p$-root pair is closely related to root systems and Weyl groups.
-In particular, we recover many results by Dolgachev–Keum with a different method and give an upper bound for the exponent of $p$ in $|G|$.
+In particular, we provide alternative proof for many results by Dolgachev–Keum and give an upper bound for the exponent of $p$ in $|G|$.
 1. Introduction
@@
 The problem of classifying (symplectic) automorphism groups of K3 surfaces over fields of positive characteristics turns out to be more complicated.
-In [DK09], Dolgachev and Keum extended Mukai’s methods to classify finite groups $G$ that act symplectically and faithfully on K3 surfaces over an algebraically closed field of characteristic $p>0$ under the tameness condition that $\gcd(|G|,p)=1$.
-They also showed that if $p\geq 13$, then such $G$ has to be tame.
-Ohashi–Schütt [OS24] recently completely classified maximal finite groups of symplectic automorphisms of superspecial K3 surfaces in any positive characteristic.
-In this paper, we investigate finite groups acting symplectically on K3 surfaces in positive characteristic $p\geq 3$ via the induced action of the group on the Néron–Severi lattice.
-Our method can be applied to group actions without the tameness assumption.
+In [DK09a], Dolgachev and Keum extended Mukai’s methods to prove that if a finite group $G$ can act symplectically and faithfully on a K3 surface over an algebraically closed field of characteristic $p>0$ coprime to $|G|$, then $G$ is isomorphic to a subgroup of $M_{23}$.
+They also showed that if $p\geq 13$, then any symplectic and faithful action of a finite group $G$ on a K3 surface in characteristic $p$ must be tame, namely, $\gcd(|G|,p)=1$.
+Using the induced action on Néron–Severi lattice, Ohashi–Schütt [OS24] recently provided a complete classification of finite groups that can act on supersingular K3 surfaces with Artin invariant $1$.
+Combining with Dogalchev–Keum [DK09a], this provides, for any prime $p$, a classification of finite groups which admit a tame symplectic action on a K3 surface in characteristic $p$.
+In this paper, we investigate finite groups acting symplectically on K3 surfaces in odd characteristic.
+Our approach is also based on the induced action on Néron–Severi lattice.
+A novelty of this work is that we uniformly deal with all K3 surfaces in odd characteristics using Leech pairs (see Definition 2.8).
+In particular, we develop a new machinery called $p$-root pairs to link the wild symplectic actions on supersingular K3 surfaces of arbitrary Artin invariants with Weyl group actions on root lattices.
+By leveraging this relation, we obtain constraints of $p$-parts of a finite group acting symplectically and faithfully on a supersingular K3 surface in odd characteristic $p$.
 Before stating our main results, we fix some terminologies.
@@
 Then we denote by $\widetilde{G}$ the group of all automorphisms of $L$ which act trivially on $A_{L}$.
-We call $\widetilde{G}$ and $(\widetilde{G},L)$ the enlargement of $G$ and $(G,L)$ respectively.
+We call $\widetilde{G}$ and $(\widetilde{G},L)$ the saturation of $G$ and $(G,L)$ respectively.
+We call $(G,L)$ saturated if $G=\widetilde{G}$.
 The K3 surface $X/k$ is said to be (Shioda-)supersingular if $\rho(X)$ reaches the maximal possibility $22$. When $X$ is supersingular,
@@
 If a K3 surface in characteristic $p$ is not supersingular, then it is of finite height with height $1\leq h\leq 10$, see discussions in §4.
-Employing the lattice-theoretic method, we will prove the following theorem.
-Theorem  1.1.
-Suppose $G$ is a finite group acting symplectically and faithfully on a K3 surface $X$ in characteristic $p\geq 3$. Suppose moreover $\gcd(|G|,p)=1$. Then
+Employing the lattice-theoretic method, we prove the following proposition.
+Proposition  1.1.
+Fix any odd prime $p$. Let $(G,S)$ be a Leech pair with $\operatorname{rank}(S)\leq 21$. The following are equivalent:
   1. (1)
-    If $X$ is supersingular with Artin invariant $\sigma$, then $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$.
-Moreover, the enlarged pair $(\widetilde{G},S_{G}(X))$ is isomorphic to a pair in Table LABEL:table:_sublist_of_HM.
+    for each supersingular K3 surface $X/k$ in characteristic $p$ with Artin invariant $\sigma$, there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
   2. (2)
-    If $X$ is of finite height $h$, then $\operatorname{rank}(S_{G}(X))\leq 21-2h$, and there exists a primitive embedding of $S_{G}(X)$ into the K3 lattice.
-In any case, $(G,S_{G}(X))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$ and $\operatorname{rank}(S_{G}(X))+\ell(A_{S_{G}(X)})\leq 22$.
+    there exists a supersingular K3 surface $X/k$ in characteristic $p$ with Artin invariant $\sigma$, such that there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
+  3. (3)
+    there exists a primitive embedding $S\hookrightarrow N_{p,\sigma}$.
+The superspecial case of Proposition 1.1 was obtained by Ohashi–Schütt [OS24].
+We collect all saturated Leech pairs $(G,S)$ that admit a primitive embedding into $N_{p,1}$ (for all odd prime $p$) in Table LABEL:table:_sublist_of_HM.
+This table was essentially obtained by Ohashi–Schütt [OS24].
+We include it in Appendix A for the reader’s convenience.
+The calculation is explained in the proof of Proposition 3.10.
+Table LABEL:table:_sublist_of_HM will be used in understanding the size of finite groups of symplectic automorphisms of supersingular K3 surfaces with arbitrary Artin invariants, see Proposition 3.12 and §6.
+An action of a finite group $G$ on a K3 surface in characteristic $p$ is called tame if $\gcd(|G|,p)=1$. We prove the following two results about tame symplectic actions.
+Proposition  1.2.
+For a tame symplectic faithful action of a finite group $G$ on a supersingular K3 surface $X$ in characteristic $p\geq 3$ with Artin invariant $\sigma$, we have $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$.
+Moreover, the saturated pair $(\widetilde{G},S_{G}(X))$ is isomorphic to a pair in Table LABEL:table:_sublist_of_HM.
+Conversely, each pair in Table LABEL:table:_sublist_of_HM (with corresponding prime $p$) is isomorphic to $(G,S_{G}(X))$ for a tame action of a finite group $G$ on a superspecial K3 surface $X$ in characteristic $p$.
+Proposition  1.3.
+For a tame symplectic faithful action of a finite group $G$ on a K3 surface $X$ of finite height $h$ in characteristic $p\geq 3$, we have $\operatorname{rank}(S_{G}(X))\leq 21-2h$, and there exists a primitive embedding of $S_{G}(X)$ into the K3 lattice.
+In both situation of Proposition 1.2 and Proposition 1.3, the pair $(G,S_{G}(X))$ is always a subpair of $(\operatorname{Co}_{0},\mathbb{L})$ and $\operatorname{rank}(S_{G}(X))+\ell(A_{S_{G}(X)})\leq 22$.
 By Höhn–Mason’s table in [HM16], a Leech pair $(G,S)$ such that $\operatorname{rank}(S)+\ell(A_{S})\leq 22$ implies that $G$ is a subgroup of $M_{23}\leq\operatorname{Co}_{0}$.
-Therefore, from Theorem 1.1, any finite group of tame symplectic automorphisms of a K3 surface in odd characteristic is a subgroup of $M_{23}$.
-This is highly compatible with the classification by Dolgachev–Keum [DK09, Theorem 4.7] and the recent result by Ohashi–Schütt [OS24, Proposition 8.2, Theorem 8.4].
+Therefore, Proposition 1.2 and Proposition 1.3 provide an alternative proof (in odd characteristic cases) for the significant result by Dolgachev–Keum [DK09a, Theorem 4.7] that any finite group of tame symplectic automorphisms of a K3 surface is a subgroup of $M_{23}$.
+See also Ohashi–Schütt [OS24, Theorem 1.1].
 Let $X$ be supersingular with Artin invariant $\sigma$.
@@
 Let $G_{R}$ be the image of $G$ in $\operatorname{Aut}(R)$, and $G_{L}=\mathrm{Ker}(G\to\operatorname{Aut}(R))$.
-Then we have exact sequence:
+Then we have an exact sequence:
 $1\rightarrow G_{L}\rightarrow G\rightarrow G_{R}\rightarrow 1$
 Further exploring the action of $G_{R}$ on $R$, we introduce the concept of (pseudo) p-root pairs (see §5).
-We summarized most of our results on supersingular K3 surfaces proved via $p$-root pairs in §5 and §6 as follows:
-Theorem  1.2.
-Let $X$ be a supersingular K3 surface over $k$, with Artin invariant $\sigma$. Let $G$ be a finite group acting faithfully and symplectically on $X$. Denote by $W(R)$ the Weyl group of $R$.
+Definition  1.4.
+Let $H$ be a finite group acting faithfully on a root lattice $R$.
+We call $(H,R)$ a pseudo $p$-root pair if the lattice $R_{H}^{\#}$ generated by $pR$ and $\{g\alpha-\alpha\big{|}g\in H,\alpha\in R\}$ contains no roots.
+We call $(H,R)$ a $p$-root pair if it is a pseudo $p$-root pair, and in addition $R^{H}=\{0\}$.
+We summarize most of our results on $p$-root pairs and their applications to supersingular K3 surfaces in §5 and §6 as follows:
+Theorem  1.5.
+Let $(H,R)$ be a $p$-root pair, then:
   1. (1)
-    If $G$ is tame then $G=G_{L}$.
+    If $R=A_{m}$, then $m+1$ is a power of $p$ (Proposition 5.6).
   2. (2)
-    $G_{R}\cap W(R)$ is always a $p$-group.
+    $R$ cannot be of type $D_{m},m\geq 5$ (Proposition 5.11). If $R=D_{4}$, $p=3$ (Subsection 5.3).
   3. (3)
-    If $p\geq 13$, $G$ is always tame.
+    If $p\geq 7$, $R$ cannot be of type $E$ (Proposition 5.20).
   4. (4)
-    If $p=11$, then the exponent of $p$ in the order of $G$ is at most $1$.
-Combining both analysis on $G_{L}$ and $G_{R}$, we also obtain an upper bound (Proposition 6.8) on the exponent of $p$ in $|G|$ for a finite group $G$ of symplectic automorphisms of a K3 surface in characteristic $p=3,5,7$.
+    $H\cap W(R)$ is always a $p$-group, here $W(R)$ denotes the Weyl group of $R$ (Proposition 5.23).
+Proposition  1.6.
+Let $X$ be a supersingular K3 surface in characteristic $p$ with Artin invariant $\sigma$. Let $G$ be a finite group acting faithfully and symplectically on $X$.
+  1. (1)
+    If $G$ is tame, then $G_{R}$ is trivial and $G=G_{L}$.
+  2. (2)
+    $(G_{R},R)$ is a $p$-root pair.
+We here point out several interesting corollaries.
+For $p\geq 13$, we deduce that a finite group acting faithfully and symplectically on a supersingular K3 surface is always tame, see Theorem 6.5.
+This provides an alternative proof of [DK09a, Theorem 2.1] for supersingular K3 surfaces.
+We also deduce that if $p=11$, then the exponent of $p$ in the order of $G$ is at most $1$. This corollary also follows from Dolgachev–Keum [DK09b, Theorem 1.1]. See Proposition 6.7.
+Moreover, combining both analysis on $G_{L}$ and $G_{R}$, we can also obtain an upper bound on the exponent of $p$ in $|G|$ for a finite group $G$ of symplectic automorphisms of a K3 surface in characteristic $p=3,5,7$.
+See Proposition 6.8.
 Now let us recall how our approach is related with earlier works.
 Kondō [Kon98] gave another proof of Mukai’s results which is lattice-theoretic and motivates our proof here.
-Ogus [Ogu79] (respectively Bragg–Lieblich [BL18]) proved the crystalline Torelli theorem for supersingular K3 surfaces in char $p\geq 5$ (respectively $p\geq 3$), which makes the lattice theory a powerful tool in the study of $\operatorname{Aut}^{s}(X)$.
-Jang [Jan16] studied the action of a symplectic automorphism on the discriminant group of the Néron–Severi group in odd characteristics which enables us to obtain Leech pairs.
+Ogus [Ogu79] (respectively Bragg–Lieblich’s unpublished work [BL18]) proved the crystalline Torelli theorem for supersingular K3 surfaces in char $p\geq 5$ (respectively $p\geq 3$), which makes the lattice theory a powerful tool in the study of $\operatorname{Aut}^{s}(X)$.
+Nygaard [Nyg81] and Jang [Jan16] studied the induced action of an automorphism on the Witt vector cohomology and the discriminant group of the Néron–Severi group in odd characteristics respectively, which enables us to obtain Leech pairs.
 The key input on lattice-theoretic side is from Höhn–Mason [HM16], Jang [Jan16] and also inspired by methods developed in Laza–Zheng [LZ22].
@@
 In particular, we give the motivation for the concept of $p$-root pairs.
-In §4, we prove Theorem 1.1 and discuss its relation with previous works.
-In §5, we developed the new machinery of (pseudo) $p$-root pairs, and in §6 we apply the new machinery to reprove an important result by Dolgachev–Keum (Theorem 1.2, (3)) in supersingular cases and give upper bounds on the wild degree of finite groups of symplectic automophisms of K3 surfaces in characteristics $p=3,5,7,11$.
+In §4, we prove Proposition 1.2, 1.3 and discuss their relation with previous works.
+In §5, we developed the new machinery of (pseudo) $p$-root pairs, and in §6 we apply the new machinery to reprove an important result by Dolgachev–Keum (Theorem 6.5) in supersingular cases and give upper bounds on the wild degree of finite groups of symplectic automophisms of K3 surfaces in characteristics $p=3,5,7,11$.
 In Appendix A, we include the full list of Höhn–Mason pairs admitting primitive embeddings in $N_{p,1}$.
+Acknowledgement: Bin Wang is supported by General Research Fund 14307022 and Early Career Scheme 24307121 of Michael McBreen.
+Zhiwei Zheng is partially supported by NSFC 12301058.
+We thank Prof.
+Looijenga for helpful discussions, especially about $p$-root pairs.
+We thank Matthias Schütt for many helpful comments after this paper being posted on arXiv.
 2. Lattice-Theoretical Results
-In this section, we introduce some terminologies and results from lattice theory that will be used. We refer to [Nik79a], [CS99], [LZ22, Appendix A] for more details.
+In this section, we introduce some terminologies and results from lattice theory that will be used. We refer to [Nik79a], [CS99, Chapter 15], [LZ22, Appendix A] for more details.
 2.1. Integral Lattices and $p$-Adic Lattices
@@
   • (I)
-    A finite symmteric bilinear form consists of a finite abelian group $A$ and a symmetric $\mathbb{Z}$-bilinear map $b:A\times A\rightarrow\mathbb{Q}/\mathbb{Z}$.
+    A finite symmetric bilinear form consists of a finite abelian group $A$ and a symmetric $\mathbb{Z}$-bilinear map $b:A\times A\rightarrow\mathbb{Q}/\mathbb{Z}$.
   • (II)
@@
 For an even integral lattice $L$, putting together the Conway–Sloane symbol of $L$ as a $p$-adic lattice for all possible $p$, we obtain the Conway–Sloane symbol of $L$.
-If two even integral lattices $L$ and $L^{\prime}$ have same Conway–Sloane symbol, then the discriminant forms $(A_{L},q_{L})$ and $(A_{L^{\prime}},q_{L^{\prime}})$ are isomorphic.
+If two even integral lattices $L$ and $L^{\prime}$ have the same Conway–Sloane symbol, then the discriminant forms $(A_{L},q_{L})$ and $(A_{L^{\prime}},q_{L^{\prime}})$ are isomorphic.
 In other words, Conway–Sloane symbol determines the discriminant form up to isomorphism.
@@
 Definition  2.8.
-For a finite group $G$ acting faithfully on a negative definite even lattice $S$, the pair $(G,S)$ is called a Leech pair, if $S$ is rootless, the induced action of $G$ on $A_{S}$ is trivial, and the action has no non-zero fixed vectors.
+[LZ22, Definition 3.3] For a finite group $G$ acting faithfully on a negative definite even lattice $S$, the pair $(G,S)$ is called a Leech pair, if $S$ is rootless, the induced action of $G$ on $A_{S}$ is trivial, and the action has no non-zero fixed vectors.
 The following criterion is crucial for us.
@@
 In particular, $G$ is a subgroup of $\operatorname{Co}_{0}$.
-This proposition is originally proved in [GHV12, appendix], see also [Huy16b, Proposition 2.2] and [LZ22, Proposition 3.4].
+This proposition is originally from [GHV12, appendix], see also [MM25, Corollary 4.19] and [Zhe25].
 2.5. Some Lemmas on Lattice Embedding
@@
 Lemma  2.11.
-Suppose a group $G$ acts on a lattice $L$, such that the induced action of $G$ on $A_{L}$ is trivial. Then $G$ acts trivially on $A_{S_{G}(L)}$.
+Suppose a group $G$ acts on an even lattice $L$, such that the induced action of $G$ on $A_{L}$ is trivial. Then $G$ acts trivially on $A_{S_{G}(L)}$.
 Proof.
@@
 2.6. Supersingular K3 Lattices
-We begin with the following proposition, in which the lattice encodes the properties of the Néron-Severi lattice of a supersingular K3 surface.
+We begin with the following proposition, in which the lattice encodes the properties of the Néron–Severi lattice of a supersingular K3 surface.
 Proposition-Definition  2.13.
@@
 $\operatorname{Pic}(X)\cong\operatorname{NS}(X)\cong\mathbb{Z}^{\oplus\rho(X)},1\leq\rho(X)\leq 22.$
-It is known that $\operatorname{NS}(X)$ is an even lattice of signature $(1,\rho(X)-1)$, see [Huy16a, §17.2]. The K3 surface $X$ is called (Shioda-)supersingular if $\rho(X)=22$.
+It is known that $\operatorname{NS}(X)$ is an even lattice of signature $(1,\rho(X)-1)$, see [Huy16, §17.2]. The K3 surface $X$ is called (Shioda-)supersingular if $\rho(X)=22$.
 Theorem  3.1.
@@
 Definition  3.2.
-[Lie13, Definition 4.14] Let $\sigma\geq 1$ be an integer and $V$ be a $2\sigma$-dimensional $\mathbb{F}_{p}$-vector space endowed with a nondegenerate and nonneutral quadratic form.
+[Ogu79, Definition 3.19][Lie13, Definition 4.14] Let $\sigma\geq 1$ be an integer and $V$ be a $2\sigma$-dimensional $\mathbb{F}_{p}$-vector space endowed with a nondegenerate and nonneutral quadratic form.
 A subspace $K\subset V\otimes_{\mathbb{F}_{p}}k$ is called characteristic if
@@
 The second map is induced by the crystalline Chern class map.
-We put $K(X,\phi)\coloneqq(id\otimes\operatorname{Fr}_{k})^{-1}(\ker(X,\phi))$.
+Ogus [Ogu79, Proposition 4.3] showed that $\ker(X,\phi)$ is a subspace of $(pN^{\vee}/pN)\otimes_{\mathbb{F}_{p}}k$.
+We then put $K(X,\phi)\coloneqq(id\otimes\operatorname{Fr}_{k})^{-1}(\ker(X,\phi))$.
 Proposition  3.3.
-[Ogu79, Proposition 4.3][Ogu83, §1.8] Both $\ker(X,\phi)$ and $K(X,\phi)$ are characteristic subspaces of $(pN^{\vee}/pN)\otimes_{\mathbb{F}_{p}}k$.
+[Ogu83, §1.8] Both $\ker(X,\phi)$ and $K(X,\phi)$ are characteristic subspaces of $(pN^{\vee}/pN)\otimes_{\mathbb{F}_{p}}k$.
 We denote the set of roots in $N$ by $\Delta$.
@@
 Jang [Jan16, Proposition 3.1] shows that for any $g\in\operatorname{Aut}(X)$, $g$ acts on $A_{\operatorname{NS}(X)}\otimes_{\mathbb{F}_{p}}k$ trivially if and only if $g$ is symplectic.
+See also [Nyg81, Theorem 2.1] and [OS24, Theorem 3.1].
 By Proposition 3.3, $K$ is contained in
@@
 We call $\mathbb{D}_{p,\sigma}$ the auxiliary lattice.
-The next proposition will be used to prove Theorem 3.8.
+The next proposition will be used to prove Proposition 1.1.
 Proposition  3.6.
@@
 ∎
-The following theorem is a criterion of Leech pairs arising from symplectic actions of finite groups on supersingular K3 surfaces:
-Theorem  3.8.
-Let $(G,S)$ be a Leech pair with $\operatorname{rank}(S)\leq 21$. The following are equivalent:
-  1. (1)
-    for each supersingular K3 surface $X/k$ with Artin invariant $\sigma$, there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
-  2. (2)
-    there exists a supersingular K3 surface $X/k$ with Artin invariant $\sigma$, such that there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
-  3. (3)
-    there exists a primitive embedding $S\hookrightarrow N_{p,\sigma}$.
-Proof.
+We now prove Proposition 1.1 which is a criterion of Leech pairs arising from symplectic actions of finite groups on supersingular K3 surfaces.
+Proof of Proposition 1.1.
 It is clear that (1) implies (2) and (2) implies (3).
@@
 Maximal finite subgroups of $\operatorname{Aut}^{s}(X)$ for superspecial K3 surfaces over an algebraically closed field of any characteristic (including $2$) are obtained recently by Ohashi–Schütt [OS24].
-For further use in this paper, we give a larger list (Table LABEL:table:_sublist_of_HM) consisting of all Höhn–Mason pairs $(G,S)$ which admit primitive embedding $S\hookrightarrow N_{p,1}$ for odd prime $p$.
-Here Höhn–Mason pair is defined as follows:
-Definition  3.9.
-A Höhn–Mason pair $(G,S)$ is a Leech pair with $G\leq\operatorname{Co}_{0}$ and $S=S_{G}(\mathbb{L})$, such that $G$ is maximal in the sense that $G$ contains all automorphisms of $S$ acting trivially on $A_{S}$, i.e., $G$ has no nontrivial enlargement.
+Together with [DK09a], they complete the classification of finite tame symplectic actions on all K3 surfaces in positive characteristic.
+We define Höhn–Mason pair as follows:
+Definition  3.8.
+A Höhn–Mason pair $(G,S)$ is a Leech pair with $G\leq\operatorname{Co}_{0}$ and $S=S_{G}(\mathbb{L})$, such that $G$ is maximal in the sense that $G$ contains all automorphisms of $S$ acting trivially on $A_{S}$, i.e., $G$ has no nontrivial saturation.
 Such kind of pairs have been classified by Höhn–Mason [HM16] and there are 290 of them.
@@
 Recall that for any Leech pair $(G,S)$, we let $\widetilde{G}$ be the group of all automorphisms of $S$ acting trivially on $A_{S}$.
-The enlargement $(\widetilde{G},S)$ is still a Leech pair.
-Proposition  3.10.
+The saturation $(\widetilde{G},S)$ is still a Leech pair.
+Proposition  3.9.
 Suppose $X$ is a superspecial K3 surface and $G\leq\operatorname{Aut}^{s}(X)$ is finite.
-Then the pair $(G,S_{G}(X))$ is a Leech pair and it can be enlarged to a Höhn–Mason pair $(\widetilde{G},S_{G}(X))$.
+Then the pair $(G,S_{G}(X))$ is a Leech pair and it can be saturated to a Höhn–Mason pair $(\widetilde{G},S_{G}(X))$.
 Proof.
@@
 ∎
-By Theorem 3.8, a Leech pair $(G,S)$ arises from a superspecial K3 surface in characteristic $p$ if and only if there exists a primitive embedding $S\hookrightarrow N_{p,1}$.
-Theorem  3.11.
+By Proposition 1.1, a Leech pair $(G,S)$ arises from a superspecial K3 surface in characteristic $p$ if and only if there exists a primitive embedding $S\hookrightarrow N_{p,1}$.
+Proposition  3.10.
 A Höhn–Mason pair $(G,S)$ admits primitive embedding $S\hookrightarrow N_{p,1}$ if and only if $(G,S,p)$ belongs to Table LABEL:table:_sublist_of_HM.
+The calculation method is similar as the proof of [LZ22, Theorem 4.14], where Laza and the second author classified symplectic automorphism groups of smooth cubic fourfolds.
 Proof.
@@
 ∎
-Remark  3.12.
-It is fine to replace the auxiliary lattice $\mathbb{D}_{p,1}$ by a negative even lattice of rank $4$ with same discriminant form.
-Remark  3.13.
-The results in Table LABEL:table:_sublist_of_HM with $\operatorname{rank}(S)=20,21$ have been obtained by Ohashi–Schütt [OS24], which is sufficient for their classification.
-The cases with $\operatorname{rank}(S)\leq 19$ will be useful to understand the size of finite groups of symplectic automorphisms of supersingular K3 surfaces with arbitrary Artin invariants, see Proposition 3.15 and §6.
-In the proof of Theorem 3.11, our description of the algorithm is brief.
-We refer to [OS24] for more details about the calculation.
-Example  3.14.
+Example  3.11.
 Take $(G,q_{S})\cong(L_{3}(4).2,4_{3}^{-1}3^{-1}7^{-1})$ for an example (the corresponding number in the main list of Höhn–Mason is 170).
@@
 Moreover we have the following characterization of $(G_{L},S_{G_{L}}(N_{p,1}))$.
-Proposition  3.15.
-The enlargement $(\widetilde{G}_{L},S_{G_{L}}(N_{p,1}))$ of $(G_{L},S_{G_{L}}(N_{p,1}))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
+Proposition  3.12.
+The saturation $(\widetilde{G}_{L},S_{G_{L}}(N_{p,1}))$ of $(G_{L},S_{G_{L}}(N_{p,1}))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
 Proof.
 This is because $S_{G_{L}}(N_{p,1})$ is a primitive sublattice of $N_{p,1}$. ∎
-Lemma  3.16.
+Lemma  3.13.
 The root lattice $R_{G}$ satisfies the following statements:
@@
 ∎
-Remark  3.17.
+Remark  3.14.
 The automorphisms of root sytems of type $A,D,E$ will come into play. Later in §5, we will systematically study root systems which admit a finite group action satisfying the conditions in Lemma 3.13.
 4. The Tame Case
-In this section we give the proof of Theorem 1.1.
-4.1. Proof of Theorem 1.1 for Supersingular K3 Surfaces
-The next lemma will be used in the proof of Theorem 1.1 for supersingular K3 surfaces.
+In this section we give the proof of Proposition 1.2 and Proposition 1.3.
+4.1. Proof of Proposition 1.2: Supersingular K3 Surfaces
+The next lemma will be useful.
 Lemma  4.1.
@@
 From Lemma 4.1 and Proposition 2.9, the Leech pair $(G,S_{G}(X))$ is isomorphic to a subpair of $(\operatorname{Co}_{0},\mathbb{L})$.
-Since $S_{G}(X)$ admits a primitive embedding into $N_{p,1}$, the enlargement of $(G,S_{G}(X))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
+Since $S_{G}(X)$ admits a primitive embedding into $N_{p,1}$, the saturation of $(G,S_{G}(X))$ is a Höhn–Mason pair belonging to Table LABEL:table:_sublist_of_HM.
 We have a saturation $S_{G}(X)\oplus\operatorname{NS}(X)^{G}\subset\operatorname{NS}(X)$.
@@
 We conclude that $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$.
-Now the proof of Theorem 1.1 in the supersingular case is complete.
-4.2. Proof of Theorem 1.1 for K3 Surfaces of Finite Height
-In this subsection, we prove Theorem 1.1 for K3 surfaces of finite height $h$.
-We mainly use Jang [Jan17] to turn this into a problem over characteristic 0.
-For more on K3 surfaces of finite height, see [AM77] or [Lie16, §6] and references therein.
+Now the proof of Proposition 1.2 is complete.
+4.2. Proof of Proposition 1.3: K3 Surfaces of Finite Height
+We mainly use Jang [Jan17] to turn this into a problem over characteristic 0. For more on K3 surfaces of finite height, see [AM77] or [Lie16, §6] and references therein.
 The following lemma is essentially due to Jang ([Jan17, Theorem 3.2, Corollary 3.4]):
@@
 Since such a conclusion is true over $\mathbb{C}$, we can use the ”Lefschetz principle” here to conclude over an algebraically closed field of characteristic 0.
-To be more precise, there exists a finitely generated field $E$ over $\mathbb{Q}$ contained in $\bar{K}$ and a K3 surfaces $X^{\prime}$ defined over $E$ with $X^{\prime}\times_{E}\overline{K}\cong\mathscr{X}_{\overline{K}}$ and $g$ acts on $X^{\prime}$ symplectically and of finite order.
+To be more precise, there exists a finitely generated field $E$ over $\mathbb{Q}$ contained in $\bar{K}$ and a K3 surface $X^{\prime}$ defined over $E$ with $X^{\prime}\times_{E}\overline{K}\cong\mathscr{X}_{\overline{K}}$ and $g$ acts on $X^{\prime}$ symplectically and of finite order.
 Since $E$ is finitely generated over $\mathbb{Q}$, there is also a field embedding $E\hookrightarrow\mathbb{C}$.
 We denote by $X^{\prime}_{\mathbb{C}}=X^{\prime}\times_{E}\mathbb{C}$.
-By [Huy16a, §17, Lemma 2.2], we have $\operatorname{NS}(\mathscr{X}_{\overline{K}})\cong\operatorname{NS}(X^{\prime}_{\overline{E}})\cong\operatorname{NS}(X^{\prime}_{\mathbb{C}})$, and the action of $g$ on $X^{\prime}_{\mathbb{C}}$ is symplectic.
+By [Huy16, §17, Lemma 2.2], we have $\operatorname{NS}(\mathscr{X}_{\overline{K}})\cong\operatorname{NS}(X^{\prime}_{\overline{E}})\cong\operatorname{NS}(X^{\prime}_{\mathbb{C}})$, and the action of $g$ on $X^{\prime}_{\mathbb{C}}$ is symplectic.
 We then know that the induced action of $g$ on $H^{2}(X^{\prime}_{\mathbb{C}},\mathbb{Z})$ fixes all elements in the transcendental lattice $T(X^{\prime}_{\mathbb{C}})$.
@@
 Notice that $\operatorname{rank}(\operatorname{NS}(X))\leq 22-2h$ (see [AM77, Page 130]). Since $G$ is finite, we have $\operatorname{rank}(S_{G}(X))\leq 21-2h$.
-Finally, since $\gcd(|G|,p)=1$, we know $\gcd(|A_{S_{G}(X)}|,p)=1$, similar as Lemma 4.1. We hence finish the proof of Theorem 1.1 for K3 surfaces of finite height.
+Finally, since $\gcd(|G|,p)=1$, we know $\gcd(|A_{S_{G}(X)}|,p)=1$, similar as Lemma 4.1. We hence finish the proof of Proposition 1.3 for K3 surfaces of finite height.
 5. $p$-Root Pairs
@@
 We will simply write $R^{\#}_{H}$ as $R^{\#}$ if no ambiguity caused.
-In the following, we will first discuss irreducible $p$-root pairs and then general $p$-root pairs. In particular, in the last subsection, we prove the second statement in Theorem 1.2.
+Suppose $R=\oplus R_{i}$, then we define $H_{i}$ to be the image of $\{g\in H\big{|}g(R_{i})=R_{i}\}$ in $\operatorname{Aut}(R_{i})$, which is called the stabilizer of $R_{i}$.
+Lemma  5.2.
+If $(H,R)$ is a pseudo $p$-root pair, then each $(H_{i},R_{i})$ is also a pseudo $p$-root pair. If $(H,R)$ is a $p$-root pair, then each $(H_{i},R_{i})$ is also a $p$-root pair.
+Proof.
+The first statement is clear.
+Suppose $(H,R)$ is a $p$-root pair, to show that $(H_{i},R_{i})$ is a $p$-root pair, we only need to show $R_{i}^{H_{i}}=\{0\}$.
+Otherwise there is a nonzero vector $v\in R_{i}^{H_{i}}$.
+The factor of $\sum\limits_{g\in H}gv$ in $R_{i}$ equals to |$|\{g\in H\big{|}g(R_{i})=R_{i}\}|v$, hence nonzero.
+Meanwhile $\sum\limits_{g\in H}gv\in R^{H}=\{0\}$, a contradiction.
+∎
+In the following, we will first discuss irreducible $p$-root pairs and then general $p$-root pairs.
+In particular, in the last subsection, we prove the last statement in Theorem 1.5.
+Actually we here show that it holds more generally for pseudo $p$-root pairs .
 5.1. $p$-Root Pairs of Type $A$
@@
 The roots $\alpha_{i}=v_{i}-v_{i+1}$ for $1\leq i\leq m$ form a base of the root lattice.
-Lemma  5.2.
+Lemma  5.3.
 Let $(H,R)$ be a pseudo $p$-root pair, with $R=A_{m}$.
@@
 ∎
-Corollary  5.3.
+Corollary  5.4.
 Let $(H,R)$ be a pseudo $p$-root pair, with $R=A_{m}$. For any $g\in H\cap W(A_{m})$, if $g$ is nontrivial, then all the orbits of $g$ have same length.
@@
 ∎
-Remark  5.4.
+Remark  5.5.
 This is a quite restrictive condition on $H$. However, $H$ needs not to be cyclic. For example,
@@
 For irreducible $p$-root pairs of type A, we have a stronger conclusion:
-Proposition  5.5.
+Proposition  5.6.
 Given a $p$-root pair $(H,R)$ with $R=A_{m}$, then $(m+1)\big{|}|H\cap W(A_{m})|$. In particular, $m+1$ is a power of $p$.
@@
 ∎
-Proposition  5.6.
+Proposition  5.7.
 If $(H,R)$ is a $p$-root pair with $R=A_{m}$, then $(H\cap W(A_{m}),R)$ is also a $p$-root pair.
@@
 ∎
-Remark  5.7.
-This may not hold in general. See Lemma 5.14.
+Remark  5.8.
+This may not hold for other root systems. See Lemma 5.15.
 5.2. $p$-Root Pairs of Type $D_{m}$ with $m\geq 5$
@@
 For any element $g\in S_{m}\ltimes(\mathbb{Z}/2\mathbb{Z})^{m}$, we denote by $\overline{g}$ its part in $S_{m}$.
-Lemma  5.8.
+Lemma  5.9.
 Suppose $(H,R)$ is a pseudo $p$-root pair of type $D_{m}$ for $m\geq 4$.
@@
 ∎
-Proposition  5.9.
+Proposition  5.10.
 Let $(H,R)$ be a pseudo $p$-root pair with $R=D_{m}$ for $m\geq 4$, then $H\cap W(R)=\{id\}$.
@@
 ∎
-Proposition  5.10.
+Proposition  5.11.
 Suppose $R=D_{m},m\geq 5$, then $(H,R)$ is a pseudo $p$-root pair if and only if:
@@
 The last statement follows if we add one more restriction that $R^{H}=\{0\}$. ∎
-Proposition  5.11.
+Proposition  5.12.
 Any $p$-root pair $(H,R)$ does not have a component of type $D_{m}$ with $m\geq 5$.
@@
 We have that $\operatorname{Aut}(D_{4})=S_{3}\ltimes W(D_{4})$.
-Consider the outer automorphisms: $x:\alpha_{3}\mapsto\alpha_{4}\mapsto\alpha_{1}\mapsto\alpha_{3},\alpha_{2}\mapsto\alpha_{2}$, $y:v_{4}\rightarrow-v_{4}$, i.e., $y:\alpha_{3}\mapsto\alpha_{4}$.
+Consider the outer automorphisms: $x:\alpha_{3}\mapsto\alpha_{4}\mapsto\alpha_{1}\mapsto\alpha_{3},\alpha_{2}\mapsto\alpha_{2}$, $y:v_{4}\rightarrow-v_{4}$, i.e., $y:\alpha_{3}\mapsto\alpha_{4}\mapsto\alpha_{3}$.
 Then the outer automorphism group $S_{3}$ is generated by $x,y$.
@@
 In other words $g=(v_{1}\mapsto-v_{1})\circ(132)\circ(v_{1}\mapsto-v_{1})$.
-Lemma  5.12.
+Lemma  5.13.
 $\langle g\rangle\times\langle x\rangle$ is a Sylow 3-subgroup of $\operatorname{Aut}(D_{4})$.
@@
 ∎
-Proposition  5.13.
+Proposition  5.14.
 If $p\geq 5$, there is no $p$-root pair of type $D_{4}$.
@@
 However, we have
-Lemma  5.14.
+Lemma  5.15.
 If $p=3$, $(\langle gx\rangle,D_{4})$ is a $p$-root pair.
@@
 […]
-If $\lambda$ is a root, then two of $\{(2a+b),(-a-b),2b,(-a)\}$ has to be 00 in $\mathbb{Z}/3\mathbb{Z}$ which is impossible. ∎
+If $\lambda$ is a root, then exactly two of $\{(2a+b),(-a-b),2b,(-a)\}$ has to be 00 in $\mathbb{Z}/3\mathbb{Z}$ which is impossible. ∎
 5.4. $p$-Root Pairs of Type $E$.
 For lattices of type $E$, we have $\operatorname{Aut}(E_{6})=W(E_{6})\rtimes(\mathbb{Z}/2\mathbb{Z})$, $\operatorname{Aut}(E_{7})=W(E_{7})$, $\operatorname{Aut}(E_{8})=W(E_{8})$.
-Lemma  5.15.
+Lemma  5.16.
 Let $(H,R)$ be a pseudo $p$-root pair, such that $|H|$ is coprime to $p$, then $S_{H}(R)$ has no roots.
@@
 ∎
-Corollary  5.16.
+Corollary  5.17.
 Suppose $(H,R)$ is a pseudo $p$-root pair, $\gcd(|H|,p)=1$ and $H\leq W(R)$. Then $(H,S_{H}(R))$ is a Leech pair.
@@
 ∎
-Proposition  5.17.
+Proposition  5.18.
 Let $(H,R)$ be a pseudo $p$-root pair and $R$ is of type $E_{n}$ where $n=6,7,8$, then $H\cap W(R)$ is a $p$-group.
@@
 ∎
-Lemma  5.18.
+Lemma  5.19.
 Suppose $(H,R)$ is a $p$-root pair with $R$ of type $E_{n}$ ($n=6,7$ or $8$), then either $p=3$ and $R=E_{6}$ or $E_{8}$, or $p=5$ and $R=E_{8}$.
@@
 For $R=E_{6}$, we first show that $H\cap W(E_{6})\neq\{id\}$.
-Otherwise $H\cap W(E_{6})=\{id\}$, then $|H|=2$, which implies $S_{H}(E_{6})$ is rootless by Lemma 5.15.
+Suppose that $H\cap W(E_{6})=\{id\}$, then $|H|=2$, which implies $S_{H}(E_{6})$ is rootless by Lemma 5.16.
 So $R^{H}$ is nonzero, contradiction.
@@
 We can conclude that:
-Proposition  5.19.
+Proposition  5.20.
 Let $(H,R)$ be a $p$-root pair. If $p\geq 7$, then $R$ has no component of type $E$.
@@
 For any root $\alpha$, we simply put $s_{\alpha}$ as the reflection defined by $\alpha$, i.e., $s_{\alpha}(x)=x-\langle x,\alpha\rangle\alpha$.
-Lemma  5.20.
+Lemma  5.21.
 There is a faithful action of $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}\subset W(A_{4})\times W(A_{4})$ on $E_{8}$.
@@
 We can verify that $ab=ba$. For example $ab(\alpha_{4})=ba(\alpha_{4})=-\theta+\sum\limits_{i=1}^{5}\alpha_{i}+\alpha_{8}$. In particular, $\langle a,b\rangle$ is a Sylow-$5$ subgroup of $W(E_{8})$. ∎
-Proposition  5.21.
+Proposition  5.22.
 If $p=5$, $(H,E_{8})$ is a $p$-root pair, then $H\cong\mathbb{Z}/5\mathbb{Z}$. Moreover, there exists a $p$-root pair $(H,E_{8})$ with $H\cong\mathbb{Z}/5\mathbb{Z}$.
@@
 5.5. On Reducible $p$-Root Pairs
-Suppose $R=\oplus R_{i}$, then we define $H_{i}$ to be the image of $\{g\in H\big{|}g(R_{i})=R_{i}\}$ in $\operatorname{Aut}(R_{i})$, which is called the stabilizer of $R_{i}$.
-Lemma  5.22.
-If $(H,R)$ is a pseudo $p$-root pair, then each $(H_{i},R_{i})$ is also a pseudo $p$-root pair. If $(H,R)$ is a $p$-root pair, then each $(H_{i},R_{i})$ is also a $p$-root pair.
-Proof.
-The first statement is clear.
-Suppose $(H,R)$ is a $p$-root pair, to show that $(H_{i},R_{i})$ is a $p$-root pair, we only need to show $R_{i}^{H_{i}}=\{0\}$.
-Otherwise there is a nonzero vector $v\in R_{i}^{H_{i}}$.
-The factor of $\sum\limits_{g\in H}gv$ in $R_{i}$ equals to |$|\{g\in H\big{|}g(R_{i})=R_{i}\}|v$, hence nonzero.
-Meanwhile $\sum\limits_{g\in H}gv\in R^{H}=\{0\}$, a contradiction.
-∎
-Lemma  5.23.
+Proposition  5.23.
 If $(H,R)$ is a pseudo $p$-root pair, then $H\cap W(R)$ is a $p$-group.
@@
 And by Lemma 3.13, $(G_{R},R)$ is a $p$-root pair.
-In this section, we apply our results on $p$-root pairs to show that when $p\geq 13$, we always have $(p,|G|)=1$, which recovers Dolgachev–Keum [DK09, Theorem 2.1] for supersingular K3.
+In this section, we apply our results on $p$-root pairs to show that when $p\geq 13$, we always have $(p,|G|)=1$.
+This provides an alternative proof of [DK09a, Theorem 2.1] for supersingular K3.
 And then we calculate the upper bounds of the $p$-part of $G$ for $p=3,5,7,11$.
@@
 ∎
-The following theorem is obtained by Dolgachev–Keum [DK09, Theorem 2.1] through a geometric argument using fixed loci and fibrations. For supersingular cases, we provide a lattice theoretic proof.
+The following theorem is obtained by Dolgachev–Keum [DK09a, Theorem 2.1] through a geometric argument using fixed loci and fibrations. For supersingular cases, we provide a lattice theoretic proof.
 Theorem  6.5 (Dolgachev–Keum).
@@
 Proof.
-For $p=11,7$, by Proposition 5.11, 5.13 and 5.19, all components of $R$ are of type A.
+For $p=11,7$, by Proposition 5.12, 5.14 and 5.20, all components of $R$ are of type A.
 Since $\operatorname{rank}R\leq 22$, by Lemma 5.2 and Proposition 5.6, each component of $R$ is isomorphic to $A_{p-1}$.
-For $p=5$, by Proposition 5.11, 5.13, there is no irreducible component of type $D$. Then the conclusion follows from Lemma 5.22, Proposition 5.5, 5.21.
-For $p=3$, by Proposition 5.11, Lemma 5.14, $D_{4}$ is the only possibility of type D. The remaining conclusion follows from Lemma 5.18, 5.22 and Proposition 5.5. ∎
+For $p=5$, by Proposition 5.12, 5.14, there is no irreducible component of type $D$. Then the conclusion follows from Lemma 5.2, Proposition 5.6, 5.22.
+For $p=3$, by Proposition 5.12, Lemma 5.15, $D_{4}$ is the only possibility of type D. The remaining conclusion follows from Lemma 5.19, 5.2 and Proposition 5.6. ∎
 Proposition  6.7.
@@
     For $p=3$, $\nu_{p}(G)\leq 14$.
+Upper bounds listed here do not depend on the Artin invariant $\sigma$.
 Proof.
@@
 For case 119, $2_{5}^{+1}4_{1}^{+1}3^{-1}5^{+1}$ is corrected to $2_{3}^{-1}4_{7}^{+1}3^{-1}5^{+1}$.
-Combining with Kondō [Kon06], in the case $(G,q_{S})\cong([2^{4}3].S_{5},8_{1}^{+1}3^{-1}5^{-1})$, we actually have $G\cong 2^{4}:(3\times S_{5})$.
+In the case $(G,q_{S})\cong([2^{4}3].S_{5},8_{1}^{+1}3^{-1}5^{-1})$, a more precise description of $G$ has been discussed in [OS24, Remark 7.2].